### v1

# LibEER: A Comprehensive Benchmark and Algorithm Library for EEG-based Emotion Recognition

Abstract
EEG-based emotion recognition (EER) is garnering increasing attention due to its potential in understanding and analyzing human emotions. Recently, significant advancements have been achieved using various deep learning-based techniques to address the EER problem. However, the absence of a convincing benchmark and open-source codebase complicates fair comparisons between different models and poses reproducibility challenges for practitioners. These issues considerably impede progress in this field. In light of this, we propose a comprehensive benchmark and algorithm library (LibEER) for fair comparisons in EER by making most of the implementation details of different methods consistent and using the same single codebase in PyTorch. In response to these challenges, we propose LibEER, a comprehensive benchmark and algorithm library for fair comparisons in EER, by ensuring consistency in the implementation details of various methods and utilizing a single codebase in PyTorch. LibEER establishes a unified evaluation framework with standardized experimental settings, enabling unbiased evaluations of over ten representative deep learning-based EER models across the four most commonly used datasets. Additionally, we conduct an exhaustive and reproducible comparison of the performance and efficiency of popular models, providing valuable insights for researchers in selecting and designing EER models. We aspire for our work to not only lower the barriers for beginners entering the field of EEG-based emotion recognition but also promote the standardization of research in this domain, thereby fostering steady development. The source code is available at https://github.com/ButterSen/LibEER.
Index Terms:
Benchmark, EEG-based emotion recognition, Fair comparison, Open source library.
I Introduction
Emotion encompasses a series of subjective cognitive experiences, serving as the primary framework for interpersonal relationships and providing the criteria for evaluating human behavior [1]. Electroencephalography (EEG) is a widely used physiological signal in emotion analysis due to its non-invasive nature, high temporal resolution, portability, and cost-effectiveness [29, 30]. Consequently, EEG-based emotion recognition (EER) has garnered increasing attention in recent years because of its robust capacity to understand emotions and has broad applicability in fields such as healthcare [31, 40], advertising [32], and education [33].
Traditional approaches primarily focus on designing or learning discriminative features, such as power spectral density (PSD) [39] and differential entropy (DE) [15], to enhance the performance of EER. Recently, with the aid of advanced neural networks, a variety of deep learning-based models [16, 18] have sprung up to recognize emotion states from EEG signals and have been one of the most promising research topics [2]. These models often aim to capture the spatiotemporal characteristics of EEG signals and can be categorized into five types: deep neural network-based (DNN), convolutional neural network-based (CNN), recurrent neural network-based (RNN), graph neural network-based (GNN), and Transformer-based methods. DNN-based methods [5, 6] are simple yet effective, employing fully connected layers and directly using the EEG signal’s features as input. CNN-based methods [8, 9, 10] treat EEG data as 2D or 3D images, where one dimension corresponds to the electrodes and the other dimension(s) represent temporal or frequency features. RNN-based methods [11], particularly the long short-term memory (LSTM) model, are suited for modeling crucial temporal information in EEG signals. GNN-based methods [12, 13, 14] are employed to model the spatial information of EEG signals across different channels. Transformer-based methods [7] are the most powerful to date, segmenting EEG signals into multiple patches and capturing dependencies between these patches using the self-attention mechanism.
Despite the significant achievements claimed by various deep learning-based models, some key issues remain underexplored in the EER research field, hindering the further development of actual progress in this area. In this paper, we identify these issues as a lack of benchmark and codebase, which makes fair comparisons difficult and leads to practitioners struggling with reproducibility.
The most critical issue pertains to the absence of a comprehensive benchmark for EER tasks, which results in non-unified, unreasonable, and ambiguous experimental settings across different models. This situation makes fair comparisons challenging and may even call into question certain conclusions. Specifically, with regard to the non-unified issue, various methods utilize differing datasets or operate under different scenarios, employ distinct pre-processing techniques (e.g., whether to conduct artifacts removal or not), feature varying ratios of training and test sets, and apply diverse data segmentation (i.e., length of samples), among other discrepancies. In terms of the unreasonable issue, certain studies employed a fixed data split strategy [15, 17, 12] -for example, utilizing the first 9 trials as training data and the final 6 trials as test data in the SEED dataset-which may result in accidental experimental outcomes. Additionally, some studies failed to satisfy cross-trial requirements in subject-dependent scenarios [34, 11], potentially leading to data leakage issues. Moreover, the majority of methods lack a validation set for model selection [13, 5], thus yielding unreliable experimental results. Regarding the ambiguous issue, many studies [7, 35, 36] do not explicitly detail how their experimental results are reported. Furthermore, data selection procedures are often unspecified, such as which session in the SEED dataset is utilized [12, 14] and how incomplete data in the HCI dataset is handled [37, 7].
An additional critical issue revolves around the absence of an open-source library for EER tasks. Despite the impressive performance demonstrated in numerous studies, the lack of publicly available code poses a significant obstacle to reproducing these outcomes [38, 5], thereby impeding researchers from building upon these studies. Furthermore, many studies fail to provide comprehensive implementation details of the models [5, 7], neglecting to include crucial information such as hyperparameters, number of network layers, activation functions, and so forth. Moreover, different studies frequently rely on distinct development platforms, such as PyTorch [19], TensorFlow [20], and MATLAB [26]. In totality, these deficiencies curtail researchers’ capacity to devise superior methods.
In this paper, we undertake the task of addressing the aforementioned critical issues and providing insights into the advancements achieved. Our approach involves the development of a comprehensive benchmark and algorithm library for EEG-based emotion recognition (referred to as LibEER). We achieve this by ensuring consistency in the implementation details of various EER methods. Firstly, we establish a fair and rational benchmark that standardizes and unifies the utilized datasets, evaluation metrics, and experimental settings. Specifically, all EER models leverage the most commonly employed datasets (SEED [22], SEED-IV [24], DEAP [21], and HCI [23]) in the field, enabling performance comparison in both subject-dependent and cross-subject scenarios. Moreover, we standardize experimental settings and methodologies, particularly by adopting a more reasonable “Train-Validation-Test” data split strategy [25, 27], widely recognized in the machine learning community for performance verification. Further, we unify the disparate length standards of EEG samples prevalent in existing research by standardizing each EEG sample length to 1s within LibEER. Additionally, we systematically implement ten representative deep learning-based EER methods with a unified training pipeline in PyTorch. This encompasses the provision of a diverse range of mainstream EEG signal preprocessing functions, evaluation protocols, data split methodologies, and experimental scenarios. To enhance the versatility of our toolkit, we furnish user-friendly interfaces for customizing experimental schemes, allowing users to easily set hyperparameters, select preprocessing methods, and modify backbone models for conducting experiments. As a result, we are able to present an accurate depiction of the current state-of-the-art results in EER.
It is noteworthy that recent research [3] has made important contributions by providing a toolbox for EER tasks. However, our work is substantially different and more incremental in several key aspects. First, we systematically analyze the problems associated with fair comparison in EER tasks, a discussion notably absent in  [3]. Second, while the existing benchmark [3] includes a variety of methods, many are not specifically tailored for EER tasks. In contrast, we have selected the ten most representative deep learning EER methods based on their impact, making our benchmarks more representative of the domain. Finally, the current benchmark [3] predominantly focuses on research with open-source code. Given that many influential EER studies do not have publicly available codes, this approach has led to the exclusion of some significant models. In response, we have carefully reproduced all representative studies including those without open-source code, which was a laborious endeavor, and performed rigorous validation of the results to ensure accuracy. We posit that our benchmark and library offer novel perspectives that can enrich the field significantly.
In summary, our main contributions are the following:
  • We have pioneered the establishment of the first benchmark specifically tailored for enabling fair comparisons in EEG-based emotion recognition. This benchmark not only unveils the actual progress of the EER task but also serves as a convenient reference for conducting comparative experiments.
  • We have introduced LibEER as an open-source project on GitHub, flexibly customizing experimental schemes with user-friendly interfaces. It facilitates a fair and comprehensive evaluation of over ten popular models across four datasets. Functioning as both a toolbox and a platform, LibEER assists practitioners in efficiently utilizing and replicating EER methods.
  • We provide comprehensive evaluations with various experimental settings, by controlling the implementation details. These experimental results offer valuable insights into the advancements made and aid researchers in the selection or design of EER models.
II Overview of EEG-based Emotion Recognition Methods
In this section, we begin by presenting the problem formulation of EEG-based emotion recognition, followed by a review of five kinds of EER methods, with an in-depth exploration of several representative approaches.
II-A Problem Formulation
Most current datasets involve recording EEG signals while participants are exposed to specific stimuli, such as videos or music. In our framework, we define a participant in an experiment as a subject. The complete set of EEG data collected from each subject under a particular stimulus condition is referred to as a trial, typically lasting between one to five minutes. Due to the high cost of acquiring EEG datasets, the number of trials within a complete dataset is generally limited. To facilitate subsequent processing tasks, the data within each trial is usually segmented; we denote this segmented data as a sample, which serves as the basic unit for further analysis. Additionally, since the EEG signals of the same subject may fluctuate at different times even when in the same emotional state, some datasets include recordings from various time periods to address the challenges of cross-temporal analysis. We refer to the data collected during a single time period as a session.
Given the definitions of subject, trial, and sample, the goal of emotion recognition is to classify the emotional state of a subject according to EEG samples. These emotional states can be dimensional, such as valence and arousal, or categorical, such as sadness and joy. In an experiment at study, given an EEG dataset $D$, we denote the subjects as $S_{i}(i=1,2,...,n)$, where $n$ refers to the number of subjects within the dataset. For each subject $S_{i}$, the trials they participated in are represented as $T^{(i)}_{j}(j=1,2,...,m)$, where $m$ is the number of trials associated with $S_{i}$. Each trial $T^{(i)}_{j}$ is further divided into samples based on the time dimension.
In subject-dependent tasks, each task is conducted using data from a single subject and the data is split based on individual trials. We denote the training, validation, and test sets for subject $S_{i}$ as $D^{(i)}_{train}$, $D^{(i)}_{val}$ and $D^{(i)}_{test}$, respectively. These sets must satisfy two essential constraints:
$D^{(i)}_{train}\cap D^{(i)}_{val}\cap D^{(i)}_{test}=\emptyset,$ (1)
$D^{(i)}_{train}\cup D^{(i)}_{val}\cup D^{(i)}_{test}=D^{(i)}.$
These conditions guarantee that there is no overlap among the trials in the training, validation, and test sets, while ensuring that the union of these sets encompasses all available data from subject $S_{i}$. Following the cross-trial principle, the $D^{(i)}_{train}$, $D^{(i)}_{val}$, $D^{(i)}_{test}$ can be defined as:
$D^{(i)}_{train}=\{T^{(i)}_{1},T^{(i)}_{2},\dots,T^{(i)}_{m_{1}}\},$ (2)
$D^{(i)}_{val}=\{T^{(i)}_{m_{1}+1},T^{(i)}_{m_{1}+2},\dots,T^{(i)}_{m_{2}}\},$
$D^{(i)}_{test}=\{T^{(i)}_{m_{2}+1},T^{(i)}_{m_{2}+2},\dots,T^{(i)}_{m}\}.$
The EEG samples from $D^{(i)}_{train}$ can be represented as $X^{(i)}_{tr}\in\mathbb{R}^{m^{(i)}_{tr}\times c\times d}$, where $m^{(i)}_{tr}$ represents the number of EEG sample of $D^{(i)}_{train}$, $c$ denotes the number of electrode channels and $d$ is the feature dimension. We denote the corresponding labels of $X^{(i)}_{tr}$ as $\boldsymbol{y}^{(i)}_{tr}\in\mathbb{R}^{m^{(i)}_{tr}}$. The subject-dependent task involves learning a mapping function $f:X^{(i)}_{tr}\rightarrow\boldsymbol{y}^{(i)}_{tr}$ with cross-entropy loss that accurately predicts the emotion state of $S_{i}$. The validation set $D^{(i)}_{val}$ is used to tune hyperparameters and select the best model, while the test set $D^{(i)}_{test}$ is used to evaluate the final model and estimate its performance on unseen data.
The cross-subject task uses EEG data from all subjects. We denote the training, validation, ans test sets as $D_{train}$, $D_{val}$, and $D_{test}$, respectively. These sets also satisfy two constraints:
$D_{train}\cap D_{val}\cap D_{test}=\emptyset,$ (3)
$D_{train}\cup D_{val}\cup D_{test}=D.$
These conditions ensure that there is no overlap subjects between three sets and the union of these sets covers all available data of the dataset. Following the cross-subject principle, the $D_{train}$, $D_{val}$, and $D_{test}$ can be defined as:
$D_{train}=\{S_{1},S_{2},\dots,S_{n_{1}}\},$ (4)
$D_{val}=\{S_{n_{1}+1},S_{n_{1}+2},\dots,S_{n_{2}}\},$
$D_{test}=\{S_{n_{2}+1},S_{n_{2}+2},\dots,S_{n}\}.$
Analogously, the EEG samples from $D_{train}$ can be represented as $X_{tr}\in\mathbb{R}^{m_{tr}\times c\times d}$, where $m_{tr}$ represents the number of EEG sample of $D_{train}$. We denote the corresponding labels of $X_{tr}$ as $\boldsymbol{y}_{tr}\in\mathbb{R}^{m_{tr}}$. The cross-subject task learns a mapping function $f:X_{tr}\rightarrow\boldsymbol{y}_{tr}$ with the cross-entropy loss. The validation set $D_{val}$ is used to tune hyperparameters and select the best model, while the test set $D_{test}$ is used to evaluate the final model and estimate its performance on unseen data.
Figure 1: The distinct utilization of EEG signals by different methods
TABLE I: Detailed Information Of Selected Baseline Methods.
Model Baseline frequency Citation count Impact factor of journal Publish time Type
DGCNN[12] 18 1105 11.2 2018 GNN
RGNN[13] 11 545 11.2 2020 GNN
EEGNet[8] 7 2979 4 2018 CNN
DBN[5] 7 1884 2.718 2015 DNN
MS-mda[6] 6 95 5.2 2021 DNN
GCBNet[14] 3 302 11.2 2019 GNN
ACRNN[11] 3 338 11.2 2020 RNN
CDCN[9] 2 142 5 2020 CNN
HSLT[7] 2 79 4.3 2022 Transformer
Tsception[10] 1 136 11.2 2022 CNN
II-B Representative Methods
We based our selection of baseline methods on a comprehensive review of recent high-quality papers in the field of EER from the past few years. After thoroughly examining 60 recent papers, we identified and analyzed 205 distinct baseline methods. Our analysis primarily focused on the frequency with which these methods were used as baselines in the surveyed papers, as this metric best reflects their authority and recognition within the field. Subsequently, we considered additional factors, such as the citation count of each method and the impact factor of the journals (or the rank of conferences) in which they were published, to refine our selection further. Lastly, we ensured a diverse range of methods by taking into account their publication time and methodological variety, including significant methods such as DNN, CNN, RNN, GNN, and Transformer. We filtered out models that were below 10% of those reported in the papers to ensure the validity of the algorithm library comparisons.
Considering the reproducibility of these methods and the limitations of our current resources, we ultimately selected the ten most representative works[8, 9, 10, 11, 12, 13, 14, 5, 6, 7] to include in LibEER. Table I presents detailed information and ranking of these ten methods. These methods represent a well-considered balance of influence, methodological diversity, and publication recency, ensuring comprehensive coverage of the most impactful and up-to-date approaches. In future work, we will continue to maintain and expand this collection by incorporating more recent and diverse methods.
II-B1 CNN-based Methods
Convolutional Neural Networks (CNN) are a type of feedforward neural network characterized by convolutional structure. When applied to EEG signals, CNN treats the EEG data as a 2D representation, where one dimension corresponds to the electrodes and the other dimension represents either temporal or frequency features of the EEG data, as shown in Figure 1 (a).
This paper discusses three key CNN-based methods. EEGNet [8] employs depthwise and separable convolutions across two blocks. The first block features temporal convolution for frequency information and depthwise convolution for spatial filters. The second block uses separable convolution to further refine the features, reducing dimensionality while preserving essential information. Channel-Fused Dense Convolutional Network (CDCN) [9] combines one-dimensional convolution with dense structures to extract temporal and spatial features from EEG signals. Its dense blocks enhance information flow and feature reuse by connecting each layer directly to all preceding layers, effectively capturing temporal dependencies and spatial correlations. TSception [10] is a multi-scale CNN that includes dynamic temporal, asymmetric spatial, and high-level fusion layers. The dynamic temporal layer captures temporal and frequency representations using multi-scale 1D kernels, while the asymmetric spatial layer learns spatial patterns. The high-level fusion layer integrates these features for efficient and accurate emotion classification.
II-B2 RNN-based Methods
Recurrent Neural Networks (RNNs) are designed for sequential data, where directed connections between nodes allow them to maintain a memory of previous inputs. In EEG applications, RNNs extract spatial features by treating channels as sequence nodes and temporal features by treating time slices as sequence nodes, effectively capturing the complex spatiotemporal dynamics in EEG signals (see Figure 1 (b)).
The Attention-based Convolutional RNN (ACRNN) [11] employs channel-wise attention and extended self-attention mechanisms. Channel-wise attention allows the model to focus on relevant channels for emotion recognition, while CNN extracts spatial features from these channels. ACRNN also uses Long Short-Term Memory (LSTM) units to capture and refine temporal features with extended self-attention, assigning importance to different time steps based on their relevance to the recognized emotional state.
II-B3 GNN-based Methods
Graph Neural Network (GNN) is a type of neural network designed to handle data structured as graphs. A graph consists of nodes and edges, which can represent complex relationships and structures. When processing EEG data, the EEG channels are typically treated as nodes in a graph, as shown in Figure 1 (c). Since EEG data does not inherently contain information about the edges or the adjacency matrix, using graph neural networks to process EEG data often involves leveraging prior knowledge about the relationships between channels or using learnable adjacency matrices as the graph’s adjacency matrix.
We introduced three GNN-based methods. Dynamical Graph CNN (DGCNN)  [12] models multichannel EEG features using a dynamic adjacency matrix that adapts during training. It employs Chebyshev graph convolution for efficient filtering calculations. Regularized GNN (RGNN)  [13] utilizes a neuroscience-inspired adjacency matrix and introduces simple graph convolution networks (SGC). RGNN enhances robustness through node-wise domain adversarial training and emotion-aware distribution learning to address cross-subject variations and noisy labels. Graph Convolutional Broad Network (GCBNet)  [14] combines graph convolutional layers with traditional convolutional layers for effective feature extraction. GCBNet uses Chebyshev graph convolution followed by stacked convolutional layers to abstract high-level features. The GCBNet+BLS variant incorporates the Broad Learning System to enhance feature extraction and classification further.
II-B4 DNN-based Methods
Deep Neural Network (DNN) is a feedforward artificial network consisting of fully connected neurons with a nonlinear activation function, including the input layer, multiple hidden layers, and the output layer. In the field of EER, DNN processes EEG data by transforming and mapping it into the feature space, enabling the extraction and utilization of complex patterns within the data, illustrated in Figure 1 (d).
In this paper, we have implemented two popular DNN-based methods. Deep Belief Networks (DBN) [5] are probabilistic generative models representing input data distributions through hidden variables. They consist of stacked restricted Boltzmann machines (RBMs), with the output of one RBM serving as input for the next, and use an efficient greedy layer-wise algorithm for pre-training. Multi-Source Marginal Distribution Adaptation (MS-MDA) [6] utilizes MLPs for feature extraction and classification while employing domain adaptation to reduce variance across subjects and sessions. MS-MDA maintains independent branches rather than merging source domains, aligning each with the target domain individually to preserve marginal distributions and enhance adaptation performance.
II-B5 Transform-based Methods
Transformer uses a self-attention mechanism to capture dependencies between elements in a sequence, effectively handling long-range dependencies. As shown in Figure 1 (e), the EEG data are typically segmented into multiple patches on channels, regions, or temporal segments when using transformers to process EEG data. These patches are then fed into the transformer model for further analysis.
The Hierarchical Spatial Learning Transformer (HSLT) [7] is a hierarchical transformer that extracts discriminative features by capturing EEG spatial dependencies from electrode to brain-region level. It inputs electrode-extracted patches into the first layer of the transformer encoder to aggregate information within the same brain region. The output then serves as input for the second layer, which captures spatial dependencies between different brain regions and highlights those with greater contributions.
TABLE II: Summary Of Our Proposed Benchmark.
Module Key Information
Datasets SEED / SEED-IV / DEAP / MAHNOB-HCI
Data Preprocessing Bandpass Filtering / Removing Eye Movement / Extracting DE features / Segmenting data
Experimental Task Subject-dependent / Cross-subject
Data Splitting Train : Val : Test = 0.6 : 0.2 : 0.2
Evaluation Methods Accuracy / F1-score
TABLE III: The Statistics of datasets in LibEER.
Dataset #Session #Subject #Trial #Channel Stimulus Duration HZ Features Labels
SEED[22] 3 15 15 62 movie clip 185-265s 200 DE positive/neutral/negative
SEED-IV[24] 3 15 24 62 movie clip 43-259s 1000 raw data happiness/sadness/fear/neutral
DEAP[21] 1 32 40 32 music video 60s 512 raw data valence/arousal
MAHNOB-HCI[23] 1 30 20 32 movie clip 34.9-117s 512 raw data valence/arousal
III Benchmark Building
A comprehensive, rational, and transparent benchmark is essential for the equitable comparison of research methodologies within the field, a need currently unmet in the EER domain. In this section, we will elaborate on the benchmark we have devised, encompassing datasets, data preprocessing, data partitioning, and evaluation methods. For convenience, we summarize the key information of the EER benchmark in Table II.
III-A Datasets
Research in the EER field exhibits variability in dataset selection and often lacks clarity regarding the specific data segments utilized. This study evaluates the performance of popular models using four commonly employed datasets, which will be introduced along with the relevant data segments, as summarized in Table III.
III-A1 SEED
The SEED[22] dataset consists of 15 subjects (7 males, 8 females), each completing three sessions, with 15 emotion-inducing video trials per session (positive, neutral, or negative, 4 minutes each). EEG data were recorded using a 62-channel cap at 200 Hz, following the 10-20 system. The dataset includes raw data and features such as PSD and DE across five frequency bands. Data from all sessions are used for subject-dependent tasks, while the first session is used for cross-subject tasks.
III-A2 SEED-IV
The SEED-IV[24] dataset comprises 15 subjects (7 males, 8 females), each participating in three sessions with 24 trials per session. Each 2-minute trial includes videos evoking happiness, sadness, fear, or neutrality. EEG data were recorded at 1000 Hz using a 62-channel cap. The dataset offers raw data and preprocessed features such as PSD and DE. All sessions’ data are used for subject-dependent tasks, while the first session is used for cross-subject tasks.
III-A3 DEAP
The DEAP[21] dataset consists of 32 subjects (16 males, 16 females) who participated in one session containing 40 trials. Each trial involved a 1-minute music video designed to evoke specific emotions, followed by participant ratings for arousal, valence, dominance, and liking on a 1-9 scale. EEG data were recorded using a 32-channel cap at 512 Hz, and the dataset includes both raw and downsampled (128 Hz) versions. All data are used for both subject-dependent and cross-subject tasks.
III-A4 MAHNOB-HCI
The MAHNOB-HCI[23] dataset includes 30 subjects (13 males, 17 females), though EEG data are completely missing for 2 subjects and partially missing for 3. Each subject participated in a single session with 20 trials, where videos lasting 34.9 to 117 seconds evoked emotions. After each video, participants rated arousal, valence, dominance, and predictability on a 1-9 scale. EEG data were recorded using a 32-channel system at 512 Hz, and the dataset includes both raw and downsampled (128 Hz) versions. Both complete and partial data are used for subject-dependent and cross-subject tasks.
III-B Data Preprocessing
A significant challenge in reproducing extant studies on EER is the lack of a standardized data preprocessing protocol. Publicly available EER datasets are not immediately usable and require extensive preprocessing. Moreover, this critical information is rarely detailed in the literature, and many open-source packages do not provide the necessary data scripts. To address this issue, our benchmark proposes a comprehensive data preprocessing framework specifically designed for EER research. The procedure includes: (1) applying bandpass filtering between 0.3 and 50 Hz; (2) eliminating eye movement artifacts using Principal Component Analysis (PCA); (3) extracting DE features across five frequency bands—[0.5, 4], [4, 8], [8, 14], [14, 30], and [30, 50]—followed by processing with a Linear Dynamic System (LDS); and (4) segmenting data using 1-second non-overlapping sliding windows to enhance the dataset.
III-C Experimental Task
The field of EER encompasses various experimental tasks, each designed to address distinct challenges with corresponding models. The primary experimental tasks are categorized as follows: (1) Subject-dependent: This task evaluates model performance for individual subjects, necessitating both training and testing data from the same subject. (2) Cross-subject: This task aims to generalize model performance across different subjects, requiring training and testing data from different subjects. (3) Subject-independent: This task assesses model performance across a group of subjects, without needing training and testing data from the same or different subjects. (4) Cross-session: This task examines model performance across different sessions of the same subject, requiring training and testing data from different sessions of the same individual.
The subject-independent task lies between the subject-dependent and cross-subject tasks, with the latter two being more aligned with real-world applications. The cross-session task necessitates datasets with multiple sessions, which many datasets lack. Therefore, our benchmark focuses on subject-dependent and cross-subject tasks, as these are the most commonly utilized and hold significant practical relevance. Additionally, it is important to note that we treat data from different sessions of the same subject as if they originate from distinct subjects.
Figure 2: The framework and data pipeline of LibEER
III-D Data Splitting
Data splitting is a critical process that delineates how a dataset is partitioned into training and test sets, significantly influencing model performance. However, prevailing practices in this domain often exhibit a lack of rationality and the absence of a coherent, standardized approach. For example, many current data-splitting methodologies fail to incorporate a validation set for model selection and do not consistently adhere to cross-trial configurations, resulting in unreliable experimental outcomes.
To mitigate these shortcomings, we propose a systematic and rational data-splitting methodology in our benchmark. Specifically, for tasks that are subject-dependent, we allocate each individual’s data into training, validation, and test sets in a 0.6:0.2:0.2 ratio, adjusting as necessary for datasets that do not divide evenly. In the case of cross-subject tasks, we similarly partition subjects into training, validation, and test sets following the same 0.6:0.2:0.2 ratio. Furthermore, our benchmark rigorously adheres to the cross-trial principle, ensuring that samples from the same trial do not appear in both the training and test sets.
III-E Evaluation Methods
The evaluation method determines how the results reported in the paper are calculated. This part of the experimental setup is also where many studies encounter the most significant issues. Currently, many studies report the best results of the model on the test set for each epoch. This evaluation method is seriously flawed as it greatly exaggerates the performance of the model.
Therefore, in our benchmark, our evaluation method reports the performance on the test set of the model that achieved the highest F1 score on the validation set across all epochs. We use accuracy and F1 score, the most commonly used metrics in existing research, and report both the mean and standard deviation. The methods for calculating accuracy and F1 score are presented in Eq. 5 and Eq. 6, respectively.
$\text{ACC}=\frac{TP+TN}{TP+TN+FP+FN}$ (5)
[…]
where TP (True Positives), TN (True Negatives), FP (False Positives), and FN (False Negatives) are the metrics used to calculate accuracy and F1 score.
IV LibEER Toolkit Framework
In this section, we will present the framework of the algorithm library to facilitate researchers’ utilization of these models in the domain of EER. Specifically, we will detail the coding procedures associated with the data loader, data split, and model training and evaluation within the LibEER. To facilitate the utilization of the LibEER library in the field of EER, users can easily install via pip, which stremlines the installation process and ensures that all necessary dependencies are managed automatically. For further details and access to the source code, please visit the official repository at https://github.com/ButterSen/LibEER.
LibEER offers a user-friendly and equitable platform for EER implemented using PyTorch. It comprises three primary modules: the data loader, data splitting, and model training and evaluation. The data loader module standardizes the data formats across various datasets and provides a range of preprocessing techniques. The data splitting module delivers a cohesive solution applicable to different experimental tasks and partitioning methods. The model training and evaluation module presents a standardized and extensible framework for conducting model training and evaluation. Figure 2 illustrates the framework of LibEER. Below, we provide detailed explanations and usage examples for each module.
Figure 3: The example shows how to LibEER to load data
Figure 4: The example how to use LibEER to slit data
IV-1 Data loader
The data loader module primarily comprises two components: dataset reading and EEG data preprocessing. Currently, the data formats in EEG emotion recognition datasets lack uniformity, necessitating distinct data loading methods for different datasets, which ultimately reduces research efficiency. To address this challenge, LibEER offers the get_uniform_data() function, which facilitates the reading of datasets through tailored methods for each dataset, subsequently integrating the data into a standardized format of (session, subject, trial). This approach not only streamlines subsequent preprocessing tasks but also enhances compatibility with various experimental tasks and data-splitting methods.
Additionally, LibEER provides the preprocess() function, which allows for the preprocessing of data based on user-defined settings. The preprocess() function operates on a trial basis and encompasses three primary steps: noise removal, feature extraction, and sample segmentation. When utilizing raw data directly from the dataset, only sample segmentation is executed. Noise removal aims to eliminate electromyography (EMG) noise, electrooculography (EOG) artifacts, and other forms of interference from the EEG signals. The feature extraction process analyzes the data to derive relevant EEG features, such as DE and PSD features, based on the specified feature types and frequency bands provided by the user. Sample segmentation organizes the data into user-defined sample lengths. Ultimately, the processed data is integrated into the format of (session, subject, trial, sample). This functionality is illustrated in the usage examples of LibEER for data loading. A data loader usage example is presented in Figure 3.
IV-2 Data split
The data splitting module is primarily responsible for partitioning the dataset into training, testing, and validation sets. LibEER performs this division based on the experimental tasks and splitting methods selected by the user. For the two experimental tasks, the merge_to_part() function is utilized to integrate the data into a standardized format.
In the context of subject-dependent tasks, the merge_to_part() function organizes the data into the format of (subject, trial, sample), wherein the data for each subject is treated as an independent sub-task, allowing for subsequent data splitting on a trial basis for each sub-task. Conversely, for cross-subject and cross-session tasks, the merge_to_part() function consolidates the data into (subject, sample) and (session, sample) formats, with data splitting conducted on a subject or session basis.
To accommodate various splitting methods, the get_split_index() function further divides the data based on the specified formats and labels. The splitting methodologies available in LibEER include fixed ratio splitting, which partitions the data into training and testing sets according to a predetermined ratio; early stopping splitting, which divides the data into training, validation, and testing sets while maintaining balanced label distribution; and cross-validation methods, which segment the data into n folds, ensuring that each fold is utilized as the testing set once. The following example (see Figure 4) illustrates how to employ LibEER for data splitting.
Figure 5: The example show how to create and train a model
TABLE IV: Experimental settings and comparison of replication results for reproduced models in LibEER.
Method Experimental Settings Results (Accuracy)
Dataset Preprocessing Task Splitting  (Train : Test) Evaluation Reported (%) Ours (%) Gap (%)
DGCNN SEED B, R, DE, 1s dependent 3 : 2 ACC 90.40$\pm$8.49 89.48$\pm$8.49 0.92$\downarrow$
RGNN SEED B, R, DE, 1s dependent 3 : 2 ACC 94.24$\pm$5.95 84.66$\pm$10.74 9.58$\downarrow$
EEGNet SEED B, R, DE, 1s dependent 3 : 2 ACC —— 68.15$\pm$12.32 ——
DBN SEED B, R, DE, 1s dependent 3 : 2 ACC 86.91$\pm$7.62 81.18$\pm$8.13 5.73$\downarrow$
MS-MDA SEED B, DE, 1s cross 14 : 1 ACC 89.63 93.97 4.34$\uparrow$
GCBNet SEED B, R, DE, 1s dependent 3 : 2 ACC 92.30$\pm$7.40 89.04$\pm$8.03 3.26$\downarrow$
GCBNet_BLS SEED B, R, DE, 1s dependent 3 : 2 ACC 94.24$\pm$6.70 88.80$\pm$9.54 5.44$\downarrow$
CDCN SEED B, R, DE, 1s dependent 3 : 2 ACC 90.63 85.10$\pm$8.80 5.53$\downarrow$
DEAP-V B, R, 1s dependent 9 : 1 ACC 92.24 92.30$\pm$11.33 0.06$\uparrow$
DEAP-A B, R, 1s dependent 9 : 1 ACC 92.92 91.99$\pm$12.20 0.93$\downarrow$
ACRNN DEAP-V B, R, DE, 3s dependent 9 : 1 ACC 93.72 86.03$\pm$9.20 7.69$\downarrow$
DEAP-A B, R, DE, 3s dependent 9 : 1 ACC 93.38 88.31$\pm$7.77 5.07$\downarrow$
HSLT DEAP-V B, R, 6s cross 14 : 1 ACC, F1 66.51 69.18 2.67$\uparrow$
DEAP-A B, R, 6s cross 14 : 1 ACC, F1 65.75 68.81 3.06$\uparrow$
DEAP-VA B, R, 6s cross 14 : 1 ACC, F1 56.93 49.57 7.36$\downarrow$
B: Bandpass filtering, R: Removing eye movement, DE: Extracting DE features, Ns: Segmenting data into N seconds.
IV-3 Model training and evaluation
Model training and evaluation are essential processes responsible for training the designated model and assessing its performance. LibEER constructs models based on the selected architecture and parameters, subsequently training them according to the specified training configurations. In scenarios where a validation set is not available, LibEER identifies the optimal performance on the testing set during training as the final outcome for that training iteration. Conversely, when a validation set is utilized, LibEER selects the network weights that yield the best performance on the validation set for evaluation against the testing set, thereby determining the final result of the training round.
The ultimate outcomes of the model will be reported according to the performance metrics specified by the user. The following example in Figure 5 illustrates the process of creating and training a model within LibEER, along with the corresponding evaluation metrics.
V Experimental Validation
In this section, we thoroughly validate the importance of LibEER and our benchmark through extensive experimentation. Initially, we utilize LibEER to replicate the selected models as closely as possible following the experimental configurations detailed in the original publications. Subsequently, we employ LibEER to compare the performance of all models based on the established benchmark. Lastly, we conduct ablation studies to address specific issues identified during the experimental phase.
TABLE V: The Mean Accuracies (and Standard Deviations) using proposed benchmark for Subject-Dependent EEG Emotion Recognition Experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
Method SVM DNN Trans CNN RNN GNN
DBN HSLT EEGNet CDCN TSception ACRNN DGCNN RGNN GCBNet GCBNet_BLS
SEED ACC 75.08 71.88 64.83 58.81 68.23 64.01 49.71 82.55 76.55 80.56 76.64
(19.73) (19.02) (20.47) (16.22) (20.35) (16.44) (13.15) (15.61) (16.92) (16.98) (17.44)
F1 70.82 67.39 58.82 54.41 63.76 60.53 45.78 79.89 72.52 77.29 72.52
(23.51) (22.81) (23.36) (17.59) (24.49) (18.51) (14.18) (18.93) (20.08) (20.92) (21.20)
SEED-IV ACC 47.80 45.56 40.28 29.89 52.26 36.06 29.01 52.39 45.40 53.28 53.51
(23.03) (21.19) (23.80) (13.53) (21.97) (15.12) (7.10) (24.32) (22.90) (21.05) (22.45)
F1 40.17 37.61 30.92 26.59 45.26 32.77 19.80 45.94 38.24 46.26 46.91
(21.68) (20.68) (24.47) (13.58) (23.00) (15.08) (5.42) (24.17) (23.09) (22.27) (22.46)
HCI-V ACC 64.83 62.03 64.00 61.15 60.48 61.12 60.51 67.83 64.86 66.84 69.60
(25.95) (24.90) (11.40) (16.76) (21.90) (15.52) (16.89) (22.40) (17.36) (21.42) (22.09)
F1 55.99 52.84 55.77 50.35 51.73 50.51 49.39 54.78 50.41 54.61 57.78
(27.80) (27.04) (13.12) (17.28) (23.11) (16.69) (15.70) (26.69) (20.34) (25.08) (27.18)
HCI-A ACC 63.61 68.51 67.74 67.42 71.82 68.26 66.26 67.29 70.96 64.89 69.60
(23.44) (21.72) (17.22) (21.71) (21.72) (23.10) (22.69) (27.73) (19.79) (27.12) (22.09)
F1 50.99 57.18 58.42 54.50 62.89 56.29 55.17 58.04 57.66 57.43 57.78
(24.89) (26.63) (19.50) (20.05) (24.99) (23.79) (23.20) (31.41) (25.39) (29.50) (27.18)
HCI-VA ACC 46.29 44.38 46.99 38.32 52.00 40.00 41.00 53.06 49.46 49.15 49.24
(28.42) (26.78) (20.76) (19.51) (26.05) (20.60) (21.58) (24.44) (23.20) (26.16) (27.97)
F1 33.91 31.96 34.76 24.56 38.04 27.19 27.10 39.75 35.97 36.49 36.68
(27.65) (27.87) (19.69) (13.71) (26.88) (13.65) (15.13) (26.01) (23.84) (27.58) (27.25)
DEAP-V ACC 54.17 56.08 56.20 51.50 57.71 51.52 53.52 56.07 55.90 56.49 57.02
(18.67) (17.38) (18.14) (11.57) (14.72) (9.54) (9.29) (17.15) (16.24) (18.17) (15.07)
F1 49.73 48.61 48.21 47.85 53.41 47.33 48.31 49.08 47.25 50.36 51.40
(18.92) (19.33) (18.73) (11.70) (15.50) (9.58) (7.77) (17.50) (17.55) (19.57) (17.25)
DEAP-A ACC 63.49 64.60 59.74 61.30 63.37 57.49 61.83 62.68 66.09 65.95 61.07
(16.72) (19.42) (18.82) (15.88) (14.18) (11.86) (14.32) (19.66) (13.91) (17.61) (16.56)
F1 53.31 52.61 50.10 53.26 53.94 50.75 49.68 53.94 49.27 55.34 50.43
(14.39) (19.85) (18.08) (13.05) (13.76) (11.30) (9.20) (20.10) (12.86) (17.86) (16.36)
DEAP-VA ACC 37.32 39.50 43.34 39.41 38.08 35.68 38.20 41.86 44.53 38.80 37.51
(17.24) (13.99) (14.49) (11.53) (15.52) (12.08) (11.34) (11.57) (14.35) (16.23) (15.29)
F1 25.55 24.88 23.47 29.19 28.90 26.75 21.05 29.12 25.88 27.89 27.00
(14.23) (10.79) (14.21) (10.21) (13.15) (8.30) (6.13) (10.92) (12.08) (15.94) (13.52)
TABLE VI: The Mean Accuracies using proposed benchmark for Cross-Subject EEG Emotion Recognition Experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
Method SVM DNN Trans CNN RNN GNN
DBN MS-MDA HSLT EEGNet CDCN TSception ACRNN DGCNN RGNN GCBNet GCBNet_BLS
SEED ACC 37.07 36.16 64.00 56.00 38.19 57.72 45.60 45.39 60.87 57.20 56.32 56.32
F1 33.45 22.67 57.35 55.75 31.83 58.66 43.54 42.37 57.22 51.37 55.12 51.43
SEED-IV ACC 28.98 36.82 56.07 30.33 28.19 31.03 34.19 31.97 42.54 44.13 32.27 40.54
F1 24.56 32.60 48.68 11.64 28.35 27.01 26.83 18.82 43.10 43.30 32.89 42.73
HCI-V ACC 70.33 69.27 67.64 66.94 57.06 67.69 57.36 54.53 63.19 65.89 65.16 71.06
F1 66.36 65.51 53.70 62.22 53.83 62.67 54.76 52.58 58.75 44.33 63.32 61.83
HCI-A ACC 54.41 57.50 57.23 49.48 54.70 55.93 52.30 51.23 59.42 57.26 52.21 60.85
F1 52.91 56.24 55.68 47.43 54.02 55.15 50.25 49.45 57.02 56.89 51.84 56.26
HCI-VA ACC 31.56 28.30 44.92 35.07 34.84 30.09 26.99 27.21 41.54 37.43 43.02 36.92
F1 26.29 27.58 21.65 27.19 27.96 24.71 21.95 20.92 38.08 19.60 36.07 32.99
DEAP-V ACC 49.58 55.02 54.82 56.56 52.36 57.78 54.44 51.94 49.91 52.15 53.58 52.34
F1 48.49 53.14 50.60 56.56 49.74 57.72 48.94 47.37 47.09 44.88 50.68 49.03
DEAP-A ACC 51.48 50.99 26.48 41.08 48.94 49.73 45.90 44.09 49.91 43.86 50.05 50.59
F1 50.95 50.09 25.39 41.05 48.94 49.17 45.56 41.51 47.09 40.46 47.79 46.38
DEAP-VA ACC 24.58 27.84 16.06 17.31 25.41 23.90 24.64 20.89 25.66 19.09 30.92 20.49
F1 24.70 24.36 15.19 16.87 24.44 22.58 23.24 15.16 24.95 13.02 30.98 18.41
V-A Model Reproduction
To ensure the reliability of our benchmark, we conducted a meticulous replication of the representative EER methods using the original configurations and compared them with the results reported in the original paper. Table IV presents both our reimplementation outcomes and those reported in the original study with the gap. Additionally, we detail the specific original experimental setups, clearly illustrating the differences in experimental details among various models.
From the results in the Table IV and the replication process, we can draw the following conclusions:
(1) The reproduced model performances are generally slightly lower than the results reported in the original paper, averaging 2.95% lower, indicating a satisfactory overall replication quality. Notably, the reproduction results for RGNN, DBN, GCBNet_BLS, and ACRNN showed more significant declines compared to the original paper. Conversely, some scenarios in MS-MDA, as well as CDCN and HSLT, exhibited improvements in replication results, while the performances of other models were comparable to those in the original study. The primary reason for these discrepancies is the lack of detailed descriptions regarding model specifications and experimental setups in some papers. We determine these parameters and settings through tuning and experience, which could introduce some biases into the results. Consequently, even when attempts are made to replicate the results as closely as possible to the descriptions in the original work, it remains challenging to achieve identical outcomes.
(2) There are significant differences in the experimental setups among various methods. The paper employs different datasets for each model, utilizes various data preprocessing techniques, executes distinct experimental tasks, and applies different data partitioning and performance evaluation methods. The substantial variability in experimental configurations prevents fair comparisons among the models. This underscores the necessity of establishing a benchmark under a unified and rational experimental framework.
V-B Model Comparison
As shown in Table IV, the implementation details of different methods vary significantly, making it difficult to fairly compare their performance and draw reliable conclusions for the advancement of the EER field. To this end, we conduct a fair comparison of all methods based on the unified and reasonable benchmark proposed in Section III. The results for the two experimental tasks are reported in Tables V and VI, respectively. We select SVM [28], one of the most commonly used machine learning algorithms, as the baseline model. Its results are presented first. In the two tables, the top two methods in each scenario are highlighted using bold and underlined formatting. From these results, we make the following key observations.
Observation 1. The graph-structured information between EEG signal channels is crucial for EER tasks. Several GNN-based models achieved superior performance.
To provide a clearer illustration of the performance of each method under the benchmark, we ranked each method’s performance in each scenario. The highest-ranking method among the $n$ methods receives $n$ points, while the lowest-ranking method receives 1 point. The total score for each method across all scenarios is then accumulated, and the scoring results for both experimental tasks are presented in Figure 6 (a) and (b).
From Figure 6 (a), it can be observed that in the subject-dependent task, DGCNN achieved the best performance with a total score of 140. Following closely are GCBNet and its variant GCBNet_BLS, both scoring 132 points. CDCN and RGNN ranked fourth and fifth with scores of 126 and 105, respectively. In contrast, EEGNet, Tsception, and ACRNN performed relatively poorly. In the cross-subject task (see Figure 6 (b)), the overall performance ranking of each model is similar to that in the subject-dependent scenario. Notably, the DBN model, which has a relatively simple architecture, rises to third place in the cross-subject scenario. In contrast, the MS-MDA model, specifically designed for cross-subject scenarios, shows significant variability across different datasets, resulting in a moderate ranking in the overall score.
In terms of model types, the four GNN-based methods demonstrated superior performance on both tasks, providing strong evidence that the graph-structured information between EEG signal channels is crucial for EER tasks. In contrast, the CNN-based methods exhibited varying levels of performance on both tasks. The DNN and Transformer-based methods showed average results on the subject-dependent task, while the DNN-based methods performed better on the cross-subject task. The RNN-based methods performed relatively poorly. However, for the Transformer and RNN-based methods, as only one model was selected from each category, the results may contain some occasionality. Based on the current results, we recommend that researchers in the EER field focus more on using GNN-based models when designing their models and emphasize the channel graph structure information inherent in EEG signals.
(a) subject-dependent
(b) cross-subject
Figure 6: The Model Performance Scoring on subject-dependent and cross-subject tasks.
(a) subject-dependent
(b) cross-subject
Figure 7: Radar chart of model performance across different datasets on subject-dependent and cross-subject tasks.
Observation 2. The issue of significant variability in EEG data among different subjects remains unaddressed, with all methods showing large standard deviations in subject-dependent scenarios and relatively low performance in cross-subject scenarios.
Table V demonstrates that the standard deviations for each method across various scenarios are significantly high, averaging approximately 15-20%. Table VI indicates that the performance of the methods is relatively low, suggesting considerable room for improvement. The pronounced variability in EEG data among subjects has long been a critical issue in the field of EER, despite some methods asserting the inclusion of specific modules to tackle this challenge. However, under our unified and systematic benchmark, the models failed to effectively address the variability in data from different subjects, as reported in the original studies. Consequently, further in-depth investigation is necessary to adequately confront the challenges posed by the substantial differences in EEG data among various subjects.
Observation 3: The scarcity of EEG data limits the effective representation learning capabilities of deep learning methods, leading to some deep learning approaches performing even worse than traditional machine learning methods, such as SVM.
As demonstrated in Tables V and VI, several carefully designed deep learning algorithms did not outperform the baseline SVM, as reported in their original studies. A significant contributing factor to this discrepancy is the scarcity of EEG data, characterized by both limited quantity and scale of datasets. Deep networks typically necessitate substantial data volumes to fully leverage their potential for representation learning. Additionally, many studies lack validation sets in their experimental designs, which can easily result in overfitting and inflated performance metrics in contexts with limited data. In contrast, our rigorous experimental setup has highlighted these shortcomings. Therefore, a prudent strategy would be to utilize various data augmentation techniques [43, 42, 41] to expand the dataset and to implement strategies aimed at reducing model overfitting.
Observation 4. The data collection methods for different EEG datasets vary, resulting in differences in data quality and task difficulty. Several methods performed better on the SEED dataset, which has a more reasonable data collection approach.
To visually illustrate the performance differences of various models across different datasets, we created radar charts for both experimental tasks, as shown in Figure 7. From the results presented in Figure 7, we summarize that models performed better on the three-class SEED dataset compared to the binary classification tasks in DEAP and HCI. We believe this is due to the higher label quality in SEED, where discrete emotion labels are easier for self-assessment, while continuous VAD labels tend to be more subjective. Additionally, SEED provides expectations for the type of emotions elicited by stimuli, whereas DEAP and HCI do not. However, as the number of labels increases, all models demonstrate similar performance on the four-class tasks of SEED-IV, DEAP-VA, and HCI-VA, with relatively poorer performance observed in DEAP-VA.
(a) Subject-dependent task on the SEED-IV dataset
(b) Cross-subject task on the SEED-IV dataset
(c) Subject-dependent task on the HCI-VA dataset
(d) Cross-subject task on the HCI-VA dataset
Figure 8: Ablation study results on sample length for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
(a) Subject-dependent task on the SEED-IV dataset
(b) Cross-subject task on the SEED-IV dataset
(c) Subject-dependent task on the HCI-VA dataset
(d) Cross-subject task on the HCI-VA dataset
Figure 9: Ablation study results on feature type for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
TABLE VII: Ablation study results on training methods for subject-dependent task on the SEED-IV and HCI-VA datasets.
Method SEED-IV HCI-VA
Ours R1 R2 Ours R1 R2
DBN ACC 45.56 49.50 43.96 44.38 46.58 42.95
(21.19) (22.02) (22.47) (26.78) (23.59) (23.24)
F1 37.61 43.34 37.83 31.96 34.87 31.87
(20.68) (22.03) (21.79) (27.87) (23.67) (21.52)
HSLT ACC 40.28 38.10 33.15 46.99 44.02 43.03
(23.80) (24.28) (20.25) (20.76) (20.09) (19.15)
F1 30.92 31.03 22.03 34.76 32.46 25.19
(24.47) (24.19) (19.80) (19.69) (18.85) (13.04)
CDCN ACC 52.26 52.18 52.83 52.00 40.41 39.84
(21.97) (21.85) (22.64) (26.05) (19.52) (23.59)
F1 45.26 44.25 47.17 38.04 24.31 25.02
(23.00) (23.53) (23.95) (26.88) (14.34) (15.63)
ACRNN ACC 29.01 27.54 27.45 41.00 41.57 38.58
(7.10) (6.86) (3.74) (21.58) (23.37) (22.25)
F1 19.80 18.79 11.74 27.10 31.78 25.34
(5.42) (6.06) (2.23) (15.13) (20.28) (14.75)
DGCNN ACC 52.39 47.89 50.01 53.06 50.16 44.38
(24.32) (23.05) (26.99) (24.44) (25.13) (21.71)
F1 45.94 41.12 43.49 39.75 37.94 31.49
(24.17) (22.44) (27.22) (26.01) (26.70) (22.08)
Average ACC 43.90 43.04 41.48 47.49 44.55 41.76
(19.68) (19.61) (19.22) (23.92) (22.34) (21.99)
F1 35.91 35.71 32.45 34.32 32.27 27.78
(19.55) (19.65) (19.00) (23.12) (20.77) (17.40)
TABLE VIII: Ablation study results on training methods for cross-subject task on the SEED-IV and HCI-VA datasets.
Method SEED-IV HCI-VA
Ours R1 R2 Ours R1 R2
DBN ACC 36.82 29.08 30.68 28.30 36.22 17.77
F1 32.60 25.52 19.79 27.58 29.53 17.56
HSLT ACC 30.33 30.33 30.33 35.07 30.77 33.12
F1 11.64 11.64 11.64 27.19 24.82 18.46
CDCN ACC 31.03 29.35 32.94 30.09 24.88 25.77
F1 27.01 27.62 31.99 24.71 20.89 21.72
ACRNN ACC 31.97 32.86 30.33 27.21 23.28 24.19
F1 18.82 20.60 11.64 20.92 17.03 16.05
DGCNN ACC 42.54 42.08 42.93 41.54 34.84 26.59
F1 43.10 40.52 42.05 38.08 33.06 20.00
Average ACC 34.54 32.74 33.44 32.44 30.00 25.49
F1 26.63 25.18 23.42 27.70 25.07 18.76
V-C Discussion of Ablation Study
In this section, we select multiple representative methods from each kind of EER method and conduct ablation studies on three crucial factors, i.e., sample length, feature type, and evaluation methods. To be specific, SEED-IV and MAHNOB-HCI with four classification labels are taken as the dataset and SVM, DBN, HSLT, CDCN, ACRNN, and DGCNN will be selected as the representative methods. The ACC and F1 scores for both experimental scenarios are reported.
V-C1 Sample length
In the ablation study on sample length, we examined lengths of 1s, 3s, 5s, 7s, and 9s, with the results presented in Figure 8. As illustrated in Figure 8, in the subject-dependent scenario, the performance of most methods tends to decline as sample length increases. We attribute this trend to the sparsity of data in the subject-dependent context; shorter sample lengths yield a greater total number of samples, which enhances the learning capacity of deep models.
In the cross-subject scenario, it can be observed from the figure that the performance of various methods with different sample lengths across the two scenarios does not show a clear pattern. Different methods may require different sample lengths in different scenarios, and choosing the appropriate length has always been a crucial and challenging task in the EER field. This suggests to researchers that the choice of sample length must balance the trade-off between the quantity of sample augmentation and the quality of the samples.
V-C2 Feature types
In the ablation study on feature types, we utilized both DE and PSD features, along with their respective variants processed by Locality Differential Sparse (LDS), as inputs. The results are presented in Figure 9. As illustrated in Figure 9, in the subject-dependent scenario, the DE and PSD features demonstrate comparable performance, while both feature types processed by LDS show significant improvements. In the cross-subject scenario, the DE feature exhibited the lowest performance, whereas the other three features displayed similar efficacy. This finding underscores the utility of LDS as a valuable feature processing technique, suggesting that the selection of specific features should be adapted flexibly according to the experimental context and the unique characteristics of the model.
V-C3 Training Methods
In the ablation study on model training methods, we aimed to explore how different training approaches impact model performance. We compared three training methods: (1) the method used in our benchmark, which selects the epoch with the best performance on the validation set (as “Ours” in Tables VII and VIII); (2) selecting the model’s result from the last epoch (as “R1” in Tables VII and VIII), where, for a fair comparison, the model’s training set includes the validation set; and (3) selecting the epoch when the loss value on the validation set decreases by less than 1% within the first 10 epochs (as “R2” in Tables VII and VIII).
Tables VII and VIII show that the training method used in our benchmark achieved the best performance in the majority of cases and obtained the highest average performance. Although “R2” also utilized the validation set like “Ours”, its performance was significantly lower than that of “Ours”. On the other hand, “R1”, despite using more data and labels in the training process, lacked a validation set, making it prone to overfitting and occasionality. Therefore, although it outperformed “Ours” in some cases, its average performance was still lower. These findings highlight the importance of using a reasonable and appropriate training method for the effective training of EER models.
VI Limitation
While LibEER is able to standardize and accelerate research in EER, we are aware of some potential limitations described as follows:
Incomplete Experimental Settings. LibEER includes several key experimental setups to facilitate research, including four widely recognized public datasets, common data preprocessing techniques, two main tasks, common data partitioning methods, and reasonable performance evaluation methods. However, LibEER does not cover all experimental configurations used in the EER field. Some studies use other datasets like DREAMER[46], Amigos[47], and MPED[48]. Others develop unique preprocessing methods, including customized filtering, artifact removal[8, 51, 52], denoising[49, 50], normalization techniques[50, 8], and extracting other types of features [12, 44, 14]. Additionally, some tasks, such as cross-session[53, 54, 55], subject-independent[54, 37], and cross-dataset tasks[45, 50], are designed to assess model performance under different conditions, such as data from different sessions or populations. Furthermore, alternative evaluation strategies, including early stopping[56, 57, 58], are employed in some studies for more robust model training and assessment. Future versions of LibEER will aim to incorporate more diverse experimental setups from the EER field.
Inadequate Model Reproduction. In replicating representative models in the EER field, we faced several challenges despite closely following the papers. Some methods lacked detailed descriptions of model or training parameters[5, 7]. In other cases [12, 13, 11], the available code did not fully align with the reported paper details. We also found that in DGCNN[12] minor adjustments during replication improved performance, suggesting the paper’s content may not have been updated. When the above issues arose, we used our experience to replicate the models as accurately as possible, adhering to the papers. The replication results (Table IV) show discrepancies compared to the original reports. In the future, we aim to collaborate with original authors and the broader EER community to improve and expand this algorithm library.
VII Conclusion
In this work, we propose LibEER, a comprehensive benchmark and algorithm library specifically designed for EEG-based emotion recognition. By standardizing the experimental settings, including dataset usage, data preprocessing methods, experimental task selection, data splitting strategies, and evaluation methods, LibEER addresses the prevalent issues of inconsistency and reproducibility in the EER field. Through its unified PyTorch codebase and the comprehensive implementation of over ten representative models, our benchmark ensures fair comparisons and accurate evaluations of models in both subject-dependent and cross-subject scenarios. Additionally, through fair comparison experiments and ablation studies, we present findings and discussions that aim to inspire researchers in the field. We believe that LibEER will not only facilitate the development and benchmarking of new EER models but also promote standardization and transparency, thereby advancing research and applications in this domain. In the future, we plan to further enhance LibEER by incorporating additional experimental settings and more advanced models.
Acknowledgments
This work was supported by National Natural Science Foundation of China (62202367), Project of China Knowledge Centre for Engineering Science and Technology.

### v2

@@
 Abstract
-EEG-based emotion recognition (EER) is garnering increasing attention due to its potential in understanding and analyzing human emotions.
-Recently, significant advancements have been achieved using various deep learning-based techniques to address the EER problem.
-However, the absence of a convincing benchmark and open-source codebase complicates fair comparisons between different models and poses reproducibility challenges for practitioners.
-These issues considerably impede progress in this field.
-In light of this, we propose a comprehensive benchmark and algorithm library (LibEER) for fair comparisons in EER by making most of the implementation details of different methods consistent and using the same single codebase in PyTorch.
-In response to these challenges, we propose LibEER, a comprehensive benchmark and algorithm library for fair comparisons in EER, by ensuring consistency in the implementation details of various methods and utilizing a single codebase in PyTorch.
-LibEER establishes a unified evaluation framework with standardized experimental settings, enabling unbiased evaluations of over ten representative deep learning-based EER models across the four most commonly used datasets.
-Additionally, we conduct an exhaustive and reproducible comparison of the performance and efficiency of popular models, providing valuable insights for researchers in selecting and designing EER models.
-We aspire for our work to not only lower the barriers for beginners entering the field of EEG-based emotion recognition but also promote the standardization of research in this domain, thereby fostering steady development.
-The source code is available at https://github.com/ButterSen/LibEER.
+EEG-based emotion recognition (EER) has gained significant attention due to its potential for understanding and analyzing human emotions.
+While recent advancements in deep learning techniques have substantially improved EER, the field lacks a convincing benchmark and comprehensive open-source libraries.
+This absence complicates fair comparisons between models and creates reproducibility challenges for practitioners, which collectively hinder progress.
+To address these issues, we introduce LibEER, a comprehensive benchmark and algorithm library designed to facilitate fair comparisons in EER.
+LibEER carefully selects popular and powerful baselines, harmonizes key implementation details across methods, and provides a standardized codebase in PyTorch.
+By offering a consistent evaluation framework with standardized experimental settings, LibEER enables unbiased assessments of over ten representative deep learning models for EER across the four most widely used datasets.
+Additionally, we conduct a thorough, reproducible comparison of model performance and efficiency, providing valuable insights to guide researchers in the selection and design of EER models.
+Moreover, we make observations and in-depth analysis on the experiment results and identify current challenges in this community.
+We hope that our work will not only lower entry barriers for newcomers to EEG-based emotion recognition but also contribute to the standardization of research in this domain, fostering steady development.
+The library and source code are publicly available at https://github.com/XJTU-EEG/LibEER.
 Index Terms:
@@
 I Introduction
-Emotion encompasses a series of subjective cognitive experiences, serving as the primary framework for interpersonal relationships and providing the criteria for evaluating human behavior [1].
-Electroencephalography (EEG) is a widely used physiological signal in emotion analysis due to its non-invasive nature, high temporal resolution, portability, and cost-effectiveness [29, 30].
-Consequently, EEG-based emotion recognition (EER) has garnered increasing attention in recent years because of its robust capacity to understand emotions and has broad applicability in fields such as healthcare [31, 40], advertising [32], and education [33].
-Traditional approaches primarily focus on designing or learning discriminative features, such as power spectral density (PSD) [39] and differential entropy (DE) [15], to enhance the performance of EER.
-Recently, with the aid of advanced neural networks, a variety of deep learning-based models [16, 18] have sprung up to recognize emotion states from EEG signals and have been one of the most promising research topics [2].
-These models often aim to capture the spatiotemporal characteristics of EEG signals and can be categorized into five types: deep neural network-based (DNN), convolutional neural network-based (CNN), recurrent neural network-based (RNN), graph neural network-based (GNN), and Transformer-based methods.
-DNN-based methods [5, 6] are simple yet effective, employing fully connected layers and directly using the EEG signal’s features as input.
-CNN-based methods [8, 9, 10] treat EEG data as 2D or 3D images, where one dimension corresponds to the electrodes and the other dimension(s) represent temporal or frequency features.
-RNN-based methods [11], particularly the long short-term memory (LSTM) model, are suited for modeling crucial temporal information in EEG signals.
-GNN-based methods [12, 13, 14] are employed to model the spatial information of EEG signals across different channels.
-Transformer-based methods [7] are the most powerful to date, segmenting EEG signals into multiple patches and capturing dependencies between these patches using the self-attention mechanism.
-Despite the significant achievements claimed by various deep learning-based models, some key issues remain underexplored in the EER research field, hindering the further development of actual progress in this area.
-In this paper, we identify these issues as a lack of benchmark and codebase, which makes fair comparisons difficult and leads to practitioners struggling with reproducibility.
-The most critical issue pertains to the absence of a comprehensive benchmark for EER tasks, which results in non-unified, unreasonable, and ambiguous experimental settings across different models.
-This situation makes fair comparisons challenging and may even call into question certain conclusions.
-Specifically, with regard to the non-unified issue, various methods utilize differing datasets or operate under different scenarios, employ distinct pre-processing techniques (e.g., whether to conduct artifacts removal or not), feature varying ratios of training and test sets, and apply diverse data segmentation (i.e., length of samples), among other discrepancies.
-In terms of the unreasonable issue, certain studies employed a fixed data split strategy [15, 17, 12] -for example, utilizing the first 9 trials as training data and the final 6 trials as test data in the SEED dataset-which may result in accidental experimental outcomes.
-Additionally, some studies failed to satisfy cross-trial requirements in subject-dependent scenarios [34, 11], potentially leading to data leakage issues.
-Moreover, the majority of methods lack a validation set for model selection [13, 5], thus yielding unreliable experimental results.
-Regarding the ambiguous issue, many studies [7, 35, 36] do not explicitly detail how their experimental results are reported.
-Furthermore, data selection procedures are often unspecified, such as which session in the SEED dataset is utilized [12, 14] and how incomplete data in the HCI dataset is handled [37, 7].
-An additional critical issue revolves around the absence of an open-source library for EER tasks.
-Despite the impressive performance demonstrated in numerous studies, the lack of publicly available code poses a significant obstacle to reproducing these outcomes [38, 5], thereby impeding researchers from building upon these studies.
-Furthermore, many studies fail to provide comprehensive implementation details of the models [5, 7], neglecting to include crucial information such as hyperparameters, number of network layers, activation functions, and so forth.
-Moreover, different studies frequently rely on distinct development platforms, such as PyTorch [19], TensorFlow [20], and MATLAB [26].
-In totality, these deficiencies curtail researchers’ capacity to devise superior methods.
-In this paper, we undertake the task of addressing the aforementioned critical issues and providing insights into the advancements achieved.
-Our approach involves the development of a comprehensive benchmark and algorithm library for EEG-based emotion recognition (referred to as LibEER).
-We achieve this by ensuring consistency in the implementation details of various EER methods.
-Firstly, we establish a fair and rational benchmark that standardizes and unifies the utilized datasets, evaluation metrics, and experimental settings.
-Specifically, all EER models leverage the most commonly employed datasets (SEED [22], SEED-IV [24], DEAP [21], and HCI [23]) in the field, enabling performance comparison in both subject-dependent and cross-subject scenarios.
-Moreover, we standardize experimental settings and methodologies, particularly by adopting a more reasonable “Train-Validation-Test” data split strategy [25, 27], widely recognized in the machine learning community for performance verification.
-Further, we unify the disparate length standards of EEG samples prevalent in existing research by standardizing each EEG sample length to 1s within LibEER.
-Additionally, we systematically implement ten representative deep learning-based EER methods with a unified training pipeline in PyTorch.
-This encompasses the provision of a diverse range of mainstream EEG signal preprocessing functions, evaluation protocols, data split methodologies, and experimental scenarios.
-To enhance the versatility of our toolkit, we furnish user-friendly interfaces for customizing experimental schemes, allowing users to easily set hyperparameters, select preprocessing methods, and modify backbone models for conducting experiments.
-As a result, we are able to present an accurate depiction of the current state-of-the-art results in EER.
-It is noteworthy that recent research [3] has made important contributions by providing a toolbox for EER tasks.
-However, our work is substantially different and more incremental in several key aspects.
-First, we systematically analyze the problems associated with fair comparison in EER tasks, a discussion notably absent in  [3].
-Second, while the existing benchmark [3] includes a variety of methods, many are not specifically tailored for EER tasks.
-In contrast, we have selected the ten most representative deep learning EER methods based on their impact, making our benchmarks more representative of the domain.
-Finally, the current benchmark [3] predominantly focuses on research with open-source code.
-Given that many influential EER studies do not have publicly available codes, this approach has led to the exclusion of some significant models.
-In response, we have carefully reproduced all representative studies including those without open-source code, which was a laborious endeavor, and performed rigorous validation of the results to ensure accuracy.
-We posit that our benchmark and library offer novel perspectives that can enrich the field significantly.
-In summary, our main contributions are the following:
-  • We have pioneered the establishment of the first benchmark specifically tailored for enabling fair comparisons in EEG-based emotion recognition.
-This benchmark not only unveils the actual progress of the EER task but also serves as a convenient reference for conducting comparative experiments.
-  • We have introduced LibEER as an open-source project on GitHub, flexibly customizing experimental schemes with user-friendly interfaces.
-It facilitates a fair and comprehensive evaluation of over ten popular models across four datasets.
-Functioning as both a toolbox and a platform, LibEER assists practitioners in efficiently utilizing and replicating EER methods.
-  • We provide comprehensive evaluations with various experimental settings, by controlling the implementation details.
-These experimental results offer valuable insights into the advancements made and aid researchers in the selection or design of EER models.
-II Overview of EEG-based Emotion Recognition Methods
-In this section, we begin by presenting the problem formulation of EEG-based emotion recognition, followed by a review of five kinds of EER methods, with an in-depth exploration of several representative approaches.
-II-A Problem Formulation
-Most current datasets involve recording EEG signals while participants are exposed to specific stimuli, such as videos or music.
-In our framework, we define a participant in an experiment as a subject.
-The complete set of EEG data collected from each subject under a particular stimulus condition is referred to as a trial, typically lasting between one to five minutes.
-Due to the high cost of acquiring EEG datasets, the number of trials within a complete dataset is generally limited.
-To facilitate subsequent processing tasks, the data within each trial is usually segmented; we denote this segmented data as a sample, which serves as the basic unit for further analysis.
-Additionally, since the EEG signals of the same subject may fluctuate at different times even when in the same emotional state, some datasets include recordings from various time periods to address the challenges of cross-temporal analysis.
-We refer to the data collected during a single time period as a session.
-Given the definitions of subject, trial, and sample, the goal of emotion recognition is to classify the emotional state of a subject according to EEG samples.
-These emotional states can be dimensional, such as valence and arousal, or categorical, such as sadness and joy.
-In an experiment at study, given an EEG dataset $D$, we denote the subjects as $S_{i}(i=1,2,...,n)$, where $n$ refers to the number of subjects within the dataset.
-For each subject $S_{i}$, the trials they participated in are represented as $T^{(i)}_{j}(j=1,2,...,m)$, where $m$ is the number of trials associated with $S_{i}$.
-Each trial $T^{(i)}_{j}$ is further divided into samples based on the time dimension.
+Emotion represents a spectrum of subjective cognitive experiences that forms the core framework for interpersonal relationships and serves as a foundation for evaluating human behaviors [1].
+Due to its non-invasive nature, high temporal resolution, portability, and cost-effectiveness, electroencephalography (EEG) has become a prominent physiological signal in emotion analysis [30, 31].
+As a result, EEG-based emotion recognition (EER) has attracted significant research interest in recent years, with applications spanning diverse fields such as healthcare [32, 41], advertising [33], and education [34].
+Traditional approaches to EER primarily rely on hand-crafted features, such as power spectral density (PSD)[40] and differential entropy (DE)[16].
+Recently, with advancements in automatic feature engineering, deep learning models have emerged as a promising research area in EER and achieved dominant performance [17, 19, 2].
+These models generally target to extract temporal features within each EEG channel, spatial features across different EEG channels, and spectral features in the frequency domain.
+Despite significant advancements in deep learning model design, the practical implementation of these approaches is often overlooked, making it difficult for newcomers and researchers to keep up with emerging studies.
+Therefore, this work focuses on the implementation aspects of EER.
+The specific issues we aim to address are as follows:
+  • I1 (Lack of benchmarking in EER): A common practice in many model design studies is to compare new approaches with state-of-the-art (SOTA) models.
+However, such comparisons are often unreliable due to the absence of comprehensive benchmarking within the EER community.
+Studies frequently use varying preprocessing techniques, such as artifact removal, and different preprocessing settings, such as varying train/test split ratios.
+Some studies also adopt questionable practices, such as trial-dependent configurations [35, 11], which can introduce data leakage risks.
+Furthermore, many studies provide ambiguous experimental settings [7, 36, 37], which hinder reproducibility.
+To address these issues, standardizing preprocessing techniques and experimental settings is essential for enabling fair comparisons and identifying genuine progress in the field.
+  • I2 (Lack of open-source libraries for deep learning-based EER): Implementation details, such as training hyperparameters, play a crucial role in model performance.
+While it is impractical and unnecessary to include every detail in an academic paper, making implementation code publicly available is vital for enabling the reproduction and further development of SOTA research.
+However, even when code is provided, it is often tied to different platforms, such as PyTorch [20], TensorFlow [21], or MATLAB [27], posing challenges to researchers.
+Therefore, there is a pressing need for an open-source library based on a standardized platform that provides comprehensive implementation details while achieving performance comparable to the original study.
+  • I3 (Lack of in-depth analysis of SOTA EER research): While deep learning has dominated EER research, many survey papers  have been published summarizing progress from various perspectives.
+However, these surveys typically report findings without hands-on experience in reproducing results, which may lead to misdirected future research.
+This lack of practical insight can obscure real challenges faced in the field.
+Therefore, an in-depth analysis of SOTA EER research based on reproduction results is essential for uncovering the true challenges and advancing the field.
+In response to these challenges, we propose LibEER (Library for EEG-based Emotion Recognition) and conduct an analysis of SOTA EER research based on reproduction results with LibEER.
+To establish a comprehensive benchmark in EER (addressing issue I1), we standardize datasets, evaluation metrics, and experimental settings.
+Specifically, we review recent SOTA EER studies to identify the most commonly used datasets, SEED [23], SEED-IV [25], DEAP [22], and MAHNOB-HCI [24], and evaluate the corresponding metrics.
+Our benchmark includes both subject-dependent and subject-independent settings, along with a variety of preprocessing methods.
+In contrast to the simplified train/test split commonly used in the literature, we adopt a more rigorous train/validation/test split to ensure a more reliable performance evaluation.
+To address issue I2, we develop the LibEER by systematically selecting and implementing ten representative deep learning-based EER methods, all within a unified training pipeline in PyTorch.
+LibEER offers a wide range of EEG signal preprocessing functions, evaluation protocols, and experimental scenarios.
+To enhance its flexibility, we provide user-friendly interfaces that allow researchers to easily customize experimental configurations.
+Users can set hyperparameters, choose preprocessing methods, and modify backbone models to suit their specific needs.
+Based on the reproduction results from LibEER, we conduct a series of comprehensive comparison experiments and identify four key observations (addressing issue I3).
+Through these experiments, we highlight both the real limitations and the progress made in current studies.
+It is worth noting that recent research [3] has made important contributions by providing a toolbox for EER.
+However, our work differs significantly and builds upon these contributions in several key areas.
+First, we systematically analyze the challenges associated with making fair comparisons in EER tasks, a discussion notably absent in [3].
+Second, while the existing benchmark [3] includes a variety of methods, many of them are not specifically tailored for EER tasks.
+In contrast, we have selected the ten most representative deep learning-based EER methods, based on their impact in this field, making our benchmark better reflect the true SOTA.
+Finally, the existing benchmark [3] predominantly focuses on research that has already made its code publicly available.
+However, many influential EER studies do not provide open-source code, which has led to the exclusion of some significant models in their work.
+In response, we have carefully reproduced all representative studies, including those without publicly available code and rigorously validated the results to ensure their accuracy.
+We believe that our benchmark and library offer fresh perspectives that will significantly advance the field.
+In summary, our main contributions are as follows:
+  • We have pioneered the establishment of the first benchmark specifically designed to enable fair comparisons in EEG-based emotion recognition (EER).
+This benchmark not only reveals the true progress of the EER task but also provides a convenient reference for conducting comparative experiments.
+  • We have introduced LibEER as an open-source project on GitHub, offering flexible customization of experimental schemes through user-friendly interfaces.
+LibEER supports fair and comprehensive evaluation of over ten popular models across four datasets.
+Serving both as a toolbox and a platform, it assists practitioners in efficiently utilizing and reproducing EER methods.
+  • We provide thorough evaluations across various experimental settings, carefully controlling implementation details.
+These experimental results offer valuable insights into the advancements made and assist researchers in selecting or designing EER models.
+The paper is organized as follows.
+Section II provides the background of the EER task.
+Section III presents the design of the benchmark.
+Section IV details the library.
+Section V covers experimental results and in-depth analysis.
+Section VI identifies future work, and Section VII concludes the whole work.
+Figure 1: Data collection protocol of EER.
+II Background of EER
+In this section, we present the general background of EEG, including data collection protocol and formal problem definition.
+II-A Data Collection Protocol
+The general protocol for EER data collection is illustrated in Fig.
+1.
+Each data collection begins with a hint of start shown to the participant.
+Afterwards, various stimuli, such as images, audio, or video, are presented to evoke specific emotions in the participant, such as happiness, sadness, or fear.
+After each stimulus, the participant is asked to self-assess their emotional state, providing a subjective emotion score based on their experience during the stimulus.
+This process is repeated a set number of times with different stimuli to collect enough data.
+In this context, the participant exposed to the stimuli, from whom data is subsequently collected, is commonly referred to as a subject.
+A continuous segment of EEG data collected from a subject under a specific stimulus condition is called a trial, typically lasting between one and five minutes.
+Between trials, the subject takes a rest.
+A series of trials is termed a session.
+Since EEG data from the same subject, even in the same emotional state, may vary significantly over time, datasets generally include several sessions, with an interval of more than 24 hours between consecutive ones.
+Given the high cost of EEG data acquisition, a trial is often divided into segments.
+Two consecutive segments within a trial can be overlapping or non-overlapping.
+Each segment is called a sample, which serves as the fundamental unit for EER.
+Notably, within a single trial, all samples should be designated for either training or test.
+A trial-dependent setting, where some samples within a trial are used for training and others for testing, is impractical for real-world applications and should be considered an invalid setup.
+II-B Problem Formulation
+Given an EER dataset $D$, a sample is denoted as $x_{i,j,k,q}\in\mathbb{R}^{c\times d}$, where $i$, $j$, $k$, and $q$ refer to the $i$-th subject, $j$-th session, $k$-th trial, and $q$-th sample within the trial, respectively.
+Here, $c$ represents the number of electrode channels, and $d$ denotes the feature dimension of each sample.
+The goal of EER is to classify a sample into a designated emotional state, which may be represented dimensionally (e.g., valence, arousal) or categorically (e.g., sadness, joy), via the following mapping:
+$\mathcal{F}:x\rightarrow\boldsymbol{y},$ (1)
+where $\boldsymbol{y}$ denotes the emotional state and $\mathcal{F}$ is a mapping function.
+SOTA map functions are usually neural networks trained on a training set $D_{\text{train}}\subset D$, and evaluated on a test set $D_{\text{test}}\subset D$.
+Additionally, a validation set $D_{\text{valid}}\subset D$ is often employed to select the best model.
+These subsets are required to be mutually exclusive, that is $D_{\text{train}}\cap D_{\text{test}}\cap D_{\text{valid}}=\emptyset$.
+TABLE I: Summary of our proposed benchmark.
+Module Key Information
+Data Preprocessing Bandpass Filtering / Removing Eye Movement / Extracting DE features / Segmenting data
+Data Splitting Train : Val : Test = 0.6 : 0.2 : 0.2
+Experimental Task Subject-dependent / Cross-subject
+Datasets SEED / SEED-IV / DEAP / MAHNOB-HCI
+Evaluation Metrics Accuracy / F1-score
+Figure 2: The distinct utilization of EEG signals by different methods
+TABLE II: Detailed information of selected baseline methods.
+Model Baseline frequency Citation count Publication Publish year Type
+Tsception[10] 2 136 TAFFC 2022 CNN
+CDCN[9] 2 142 TCDS 2020 CNN
+EEGNet[8] 7 2979 JNE 2018 CNN
+R2G-STNN[14] 8 235 TAFFC 2022 RNN
+BiDANN[36] 4 250 TAFFC 2021 RNN
+ACRNN[11] 3 338 TAFFC 2020 RNN
+MS-mda[6] 6 95 Front. Neurosci. 2021 DNN
+DBN[5] 7 1884 TAMD 2015 DNN
+RGNN[13] 11 545 TAFFC 2020 GNN
+GCBNet[15] 3 302 TAFFC 2019 GNN
+DGCNN[12] 18 1105 TAFFC 2018 GNN
+HSLT[7] 2 79 IEEE SENS J 2022 Transformer
+III Benchmark Building
+A comprehensive, well-structured, and transparent benchmark is essential for fair comparisons of research methodologies in the field—a need that remains unmet in the EER domain.
+In this section, we present our benchmark, covering baseline selection, preprocessing, experimental setups, datasets, and evaluation protocols.
+For ease of reference, TABLE I provides a summary of the key information in our EER benchmark.
+III-A Baseline Selection
+We select representative models for the benchmark. Specifically, we set up the following criteria:
+  • We first collected a total of 60 high-quality papers published in the past year in the field of EEG-based Emotion Recognition (EER) from reputable journals and conferences (such as TAFFC, JBHI, and ICASSP).
+  • We then compiled a list of 205 comparative methods cited across these 60 papers.
+  • After that, we ranked these 205 methods based on their frequency as comparison methods in the 60 papers and selected the top 20 for replication.
+  • These 20 methods, based on model architecture, are divided into five categories: CNN, RNN, GNN, DNN, and Transformer.
+Finally, we selected models with replication results within a 10% error margin compared to the original reported results, while ensuring that each category includes at least one representative model.
+Based on these steps, we ultimately identified the twelve most representative works [8, 9, 10, 11, 12, 13, 15, 5, 6, 7, 36, 14] to include in LibEER.
+TABLE II presents detailed information on these twelve methods.
+In future work, we will continue to maintain and expand this collection by incorporating more recent and diverse methods.
+The following subsections will offer detailed descriptions of the five categories, as well as an examination of each selected method.
+III-A1 CNN-based Methods
+Convolutional Neural Networks (CNNs) are a type of feedforward neural network characterized by their convolutional operations.
+When applied to EEG signals, CNNs usually treat EEG data as a two-dimensional (2D) representation: one dimension corresponds to the channels, while the other dimension represents either the temporal or spectral features of the EEG data.
+This structure is illustrated in Fig.
+2 (a), which could handle temporal/spectral and spatial information simultaneously.
+On the other hand, some work relies on 1D CNN to extract temporal and spatial features separately.
+LibEER includes three representative CNN-based methods: EEGNet [8], CDCN [9], and TSception [10].
+Although not initially proposed for EER, EEGNet is included as a representative 2D CNN-based model due to its significant influence, with over 3000 citations.
+Both CDCN and TSception employ 1D CNNs for spatial and temporal feature extraction.
+TSception, a more advanced approach, incorporates multi-scale 1D CNNs for temporal features and employs hemisphere kernels for spatial features.
+III-A2 RNN-based Methods
+Recurrent Neural Networks (RNNs) are inherently designed for sequential data, making them well-suited for extracting temporal features from EEG signals.
+For spatial feature extraction, RNNs treat channels as sequence nodes, as illustrated in Fig.
+2 (b).
+In practice, Long Short-Term Memory (LSTM) is a widely used variant of RNNs.
+Our benchmark includes three RNN-based methods: ACRNN [11], BiDANN [36], and R2G-STNN [14].
+Both ACRNN and BiDANN use LSTM for temporal feature extraction, while R2G-STNN applies bidirectional LSTM to capture spatial features.
+An attention mechanism is often integrated with LSTM to enhance feature extraction; for instance, ACRNN uses channel-wise attention for spatial features and extended self-attention with LSTM for temporal feature extraction.
+III-A3 GNN-based Methods
+Graph Neural Network (GNN) is a type of neural network designed to handle data structured as graphs.
+A graph consists of nodes and edges, which can represent complex relationships and structures.
+When processing EEG data, the EEG channels are typically treated as nodes in a graph, as shown in Fig.
+2 (c).
+Since EEG data does not inherently contain information about the edges or the adjacency matrix, using graph neural networks to process EEG data often involves leveraging prior knowledge about the relationships between channels or using learnable adjacency matrices as the graph’s adjacency matrix.
+We introduced three GNN-based methods.
+Dynamical Graph CNN (DGCNN)  [12] models multichannel EEG features using a dynamic adjacency matrix that adapts during training.
+It employs Chebyshev graph convolution for efficient filtering calculations.
+Regularized GNN (RGNN)  [13] utilizes a neuroscience-inspired adjacency matrix and introduces simple graph convolution networks (SGC).
+RGNN enhances robustness through node-wise domain adversarial training and emotion-aware distribution learning to address cross-subject variations and noisy labels.
+Graph Convolutional Broad Network (GCBNet)  [15] combines graph convolutional layers with traditional convolutional layers for effective feature extraction.
+GCBNet uses Chebyshev graph convolution followed by stacked convolutional layers to abstract high-level features.
+The GCBNet+BLS variant incorporates the Broad Learning System to enhance feature extraction and classification further.
+III-A4 DNN-based Methods
+Deep Neural Network (DNN) is a feedforward artificial network consisting of fully connected neurons with a nonlinear activation function, including the input layer, multiple hidden layers, and the output layer.
+In the field of EER, DNN processes EEG data by transforming and mapping it into the feature space, enabling the extraction and utilization of complex patterns within the data, illustrated in Fig.
+2 (d).
+In this paper, we have implemented two popular DNN-based methods.
+Deep Belief Networks (DBN) [5] are probabilistic generative models representing input data distributions through hidden variables.
+They consist of stacked restricted Boltzmann machines (RBMs), with the output of one RBM serving as input for the next, and use an efficient greedy layer-wise algorithm for pre-training.
+Multi-Source Marginal Distribution Adaptation (MS-MDA) [6] utilizes MLPs for feature extraction and classification while employing domain adaptation to reduce variance across subjects and sessions.
+MS-MDA maintains independent branches rather than merging source domains, aligning each with the target domain individually to preserve marginal distributions and enhance adaptation performance.
+III-A5 Transform-based Methods
+Transformer uses a self-attention mechanism to capture dependencies between elements in a sequence, effectively handling long-range dependencies.
+As shown in Fig.
+2 (e), the EEG data are typically segmented into multiple patches on channels, regions, or temporal segments when using transformers to process EEG data.
+These patches are then fed into the transformer model for further analysis.
+The Hierarchical Spatial Learning Transformer (HSLT) [7] is a hierarchical transformer that extracts discriminative features by capturing EEG spatial dependencies from electrode to brain-region level.
+It inputs electrode-extracted patches into the first layer of the transformer encoder to aggregate information within the same brain region.
+The output then serves as input for the second layer, which captures spatial dependencies between different brain regions and highlights those with greater contributions.
+III-B Data preprocessing and Splitting
+III-B1 preprocessing
+A significant challenge in reproducing extant studies on EER is the lack of a standardized data preprocessing protocol.
+Publicly available EER datasets are not immediately usable and require extensive preprocessing.
+Moreover, this critical information is rarely detailed in the literature, and many open-source packages do not provide the necessary data scripts.
+To address this issue, our benchmark proposes a comprehensive data preprocessing framework specifically designed for EER research.
+The procedure includes: (1) applying bandpass filtering between 0.3 and 50 Hz; (2) eliminating eye movement artifacts using Principal Component Analysis (PCA); (3) extracting DE features across five frequency bands—[0.5, 4], [4, 8], [8, 14], [14, 30], and [30, 50]—followed by processing with a Linear Dynamic System (LDS); and (4) segmenting data using 1-second non-overlapping sliding windows to enhance the dataset.
+III-B2 Splitting
+Data splitting is a critical process that delineates how a dataset is partitioned into training and test sets, significantly influencing model performance.
+However, prevailing practices in this domain often exhibit a lack of rationality and the absence of a coherent, standardized approach.
+For example, many current data-splitting methodologies fail to incorporate a validation set for model selection and do not consistently adhere to cross-trial configurations, resulting in unreliable experimental outcomes.
+To mitigate these shortcomings, we propose a systematic and rational data-splitting methodology in our benchmark.
+Specifically, for tasks that are subject-dependent, we allocate each individual’s data into training, validation, and test sets in a 0.6:0.2:0.2 ratio, adjusting as necessary for datasets that do not divide evenly.
+In the case of cross-subject tasks, we similarly partition subjects into training, validation, and test sets following the same 0.6:0.2:0.2 ratio.
+Furthermore, our benchmark rigorously adheres to the cross-trial principle, ensuring that samples from the same trial do not appear in both the training and test sets.
+III-C Experimental Tasks
+The field of EER encompasses various experimental tasks, each designed to address distinct challenges with corresponding models.
+The primary experimental tasks are categorized as follows: (1) Subject-dependent: This task evaluates model performance for individual subjects, necessitating both training and testing data from the same subject.
+(2) Cross-subject: This task aims to generalize model performance across different subjects, requiring training and testing data from different subjects.
+(3) Subject-independent: This task assesses model performance across a group of subjects, without needing training and testing data from the same or different subjects.
+(4) Cross-session: This task examines model performance across different sessions of the same subject, requiring training and testing data from different sessions of the same individual.
+The subject-independent task lies between the subject-dependent and cross-subject tasks, with the latter two being more aligned with real-world applications.
+The cross-session task necessitates datasets with multiple sessions, which many datasets lack.
+Therefore, our benchmark focuses on subject-dependent and cross-subject tasks, as these are the most commonly utilized and hold significant practical relevance.
+Additionally, it is important to note that we treat data from different sessions of the same subject as if they originate from distinct subjects.
+We provide specific definitions for the subject-dependent and cross-subject tasks below.
 In subject-dependent tasks, each task is conducted using data from a single subject and the data is split based on individual trials.
@@
 The validation set $D_{val}$ is used to tune hyperparameters and select the best model, while the test set $D_{test}$ is used to evaluate the final model and estimate its performance on unseen data.
-Figure 1: The distinct utilization of EEG signals by different methods
-TABLE I: Detailed Information Of Selected Baseline Methods.
-Model Baseline frequency Citation count Impact factor of journal Publish time Type
-DGCNN[12] 18 1105 11.2 2018 GNN
-RGNN[13] 11 545 11.2 2020 GNN
-EEGNet[8] 7 2979 4 2018 CNN
-DBN[5] 7 1884 2.718 2015 DNN
-MS-mda[6] 6 95 5.2 2021 DNN
-GCBNet[14] 3 302 11.2 2019 GNN
-ACRNN[11] 3 338 11.2 2020 RNN
-CDCN[9] 2 142 5 2020 CNN
-HSLT[7] 2 79 4.3 2022 Transformer
-Tsception[10] 1 136 11.2 2022 CNN
-II-B Representative Methods
-We based our selection of baseline methods on a comprehensive review of recent high-quality papers in the field of EER from the past few years.
-After thoroughly examining 60 recent papers, we identified and analyzed 205 distinct baseline methods.
-Our analysis primarily focused on the frequency with which these methods were used as baselines in the surveyed papers, as this metric best reflects their authority and recognition within the field.
-Subsequently, we considered additional factors, such as the citation count of each method and the impact factor of the journals (or the rank of conferences) in which they were published, to refine our selection further.
-Lastly, we ensured a diverse range of methods by taking into account their publication time and methodological variety, including significant methods such as DNN, CNN, RNN, GNN, and Transformer.
-We filtered out models that were below 10% of those reported in the papers to ensure the validity of the algorithm library comparisons.
-Considering the reproducibility of these methods and the limitations of our current resources, we ultimately selected the ten most representative works[8, 9, 10, 11, 12, 13, 14, 5, 6, 7] to include in LibEER.
-Table I presents detailed information and ranking of these ten methods.
-These methods represent a well-considered balance of influence, methodological diversity, and publication recency, ensuring comprehensive coverage of the most impactful and up-to-date approaches.
-In future work, we will continue to maintain and expand this collection by incorporating more recent and diverse methods.
-II-B1 CNN-based Methods
-Convolutional Neural Networks (CNN) are a type of feedforward neural network characterized by convolutional structure.
-When applied to EEG signals, CNN treats the EEG data as a 2D representation, where one dimension corresponds to the electrodes and the other dimension represents either temporal or frequency features of the EEG data, as shown in Figure 1 (a).
-This paper discusses three key CNN-based methods.
-EEGNet [8] employs depthwise and separable convolutions across two blocks.
-The first block features temporal convolution for frequency information and depthwise convolution for spatial filters.
-The second block uses separable convolution to further refine the features, reducing dimensionality while preserving essential information.
-Channel-Fused Dense Convolutional Network (CDCN) [9] combines one-dimensional convolution with dense structures to extract temporal and spatial features from EEG signals.
-Its dense blocks enhance information flow and feature reuse by connecting each layer directly to all preceding layers, effectively capturing temporal dependencies and spatial correlations.
-TSception [10] is a multi-scale CNN that includes dynamic temporal, asymmetric spatial, and high-level fusion layers.
-The dynamic temporal layer captures temporal and frequency representations using multi-scale 1D kernels, while the asymmetric spatial layer learns spatial patterns.
-The high-level fusion layer integrates these features for efficient and accurate emotion classification.
-II-B2 RNN-based Methods
-Recurrent Neural Networks (RNNs) are designed for sequential data, where directed connections between nodes allow them to maintain a memory of previous inputs.
-In EEG applications, RNNs extract spatial features by treating channels as sequence nodes and temporal features by treating time slices as sequence nodes, effectively capturing the complex spatiotemporal dynamics in EEG signals (see Figure 1 (b)).
-The Attention-based Convolutional RNN (ACRNN) [11] employs channel-wise attention and extended self-attention mechanisms.
-Channel-wise attention allows the model to focus on relevant channels for emotion recognition, while CNN extracts spatial features from these channels.
-ACRNN also uses Long Short-Term Memory (LSTM) units to capture and refine temporal features with extended self-attention, assigning importance to different time steps based on their relevance to the recognized emotional state.
-II-B3 GNN-based Methods
-Graph Neural Network (GNN) is a type of neural network designed to handle data structured as graphs.
-A graph consists of nodes and edges, which can represent complex relationships and structures.
-When processing EEG data, the EEG channels are typically treated as nodes in a graph, as shown in Figure 1 (c).
-Since EEG data does not inherently contain information about the edges or the adjacency matrix, using graph neural networks to process EEG data often involves leveraging prior knowledge about the relationships between channels or using learnable adjacency matrices as the graph’s adjacency matrix.
-We introduced three GNN-based methods.
-Dynamical Graph CNN (DGCNN)  [12] models multichannel EEG features using a dynamic adjacency matrix that adapts during training.
-It employs Chebyshev graph convolution for efficient filtering calculations.
-Regularized GNN (RGNN)  [13] utilizes a neuroscience-inspired adjacency matrix and introduces simple graph convolution networks (SGC).
-RGNN enhances robustness through node-wise domain adversarial training and emotion-aware distribution learning to address cross-subject variations and noisy labels.
-Graph Convolutional Broad Network (GCBNet)  [14] combines graph convolutional layers with traditional convolutional layers for effective feature extraction.
-GCBNet uses Chebyshev graph convolution followed by stacked convolutional layers to abstract high-level features.
-The GCBNet+BLS variant incorporates the Broad Learning System to enhance feature extraction and classification further.
-II-B4 DNN-based Methods
-Deep Neural Network (DNN) is a feedforward artificial network consisting of fully connected neurons with a nonlinear activation function, including the input layer, multiple hidden layers, and the output layer.
-In the field of EER, DNN processes EEG data by transforming and mapping it into the feature space, enabling the extraction and utilization of complex patterns within the data, illustrated in Figure 1 (d).
-In this paper, we have implemented two popular DNN-based methods.
-Deep Belief Networks (DBN) [5] are probabilistic generative models representing input data distributions through hidden variables.
-They consist of stacked restricted Boltzmann machines (RBMs), with the output of one RBM serving as input for the next, and use an efficient greedy layer-wise algorithm for pre-training.
-Multi-Source Marginal Distribution Adaptation (MS-MDA) [6] utilizes MLPs for feature extraction and classification while employing domain adaptation to reduce variance across subjects and sessions.
-MS-MDA maintains independent branches rather than merging source domains, aligning each with the target domain individually to preserve marginal distributions and enhance adaptation performance.
-II-B5 Transform-based Methods
-Transformer uses a self-attention mechanism to capture dependencies between elements in a sequence, effectively handling long-range dependencies.
-As shown in Figure 1 (e), the EEG data are typically segmented into multiple patches on channels, regions, or temporal segments when using transformers to process EEG data.
-These patches are then fed into the transformer model for further analysis.
-The Hierarchical Spatial Learning Transformer (HSLT) [7] is a hierarchical transformer that extracts discriminative features by capturing EEG spatial dependencies from electrode to brain-region level.
-It inputs electrode-extracted patches into the first layer of the transformer encoder to aggregate information within the same brain region.
-The output then serves as input for the second layer, which captures spatial dependencies between different brain regions and highlights those with greater contributions.
-TABLE II: Summary Of Our Proposed Benchmark.
-Module Key Information
-Datasets SEED / SEED-IV / DEAP / MAHNOB-HCI
-Data Preprocessing Bandpass Filtering / Removing Eye Movement / Extracting DE features / Segmenting data
-Experimental Task Subject-dependent / Cross-subject
-Data Splitting Train : Val : Test = 0.6 : 0.2 : 0.2
-Evaluation Methods Accuracy / F1-score
 TABLE III: The Statistics of datasets in LibEER.
@@
 MAHNOB-HCI[24] 1 30 20 32 movie clip 34.9-117s 512 raw data valence/arousal
-III Benchmark Building
-A comprehensive, rational, and transparent benchmark is essential for the equitable comparison of research methodologies within the field, a need currently unmet in the EER domain.
-In this section, we will elaborate on the benchmark we have devised, encompassing datasets, data preprocessing, data partitioning, and evaluation methods.
-For convenience, we summarize the key information of the EER benchmark in Table II.
-III-A Datasets
+Figure 3: The framework and data pipeline of LibEER
+III-D Datasets
 Research in the EER field exhibits variability in dataset selection and often lacks clarity regarding the specific data segments utilized.
-This study evaluates the performance of popular models using four commonly employed datasets, which will be introduced along with the relevant data segments, as summarized in Table III.
-III-A1 SEED
+This study evaluates the performance of popular models using four datasets, which are selected by comprehensively considering factors such as their frequency of use, data scale, and data quality.
+They will be introduced along with the relevant data segments, as summarized in Table III.
+III-D1 SEED
 The SEED[23] dataset consists of 15 subjects (7 males, 8 females), each completing three sessions, with 15 emotion-inducing video trials per session (positive, neutral, or negative, 4 minutes each).
@@
 Data from all sessions are used for subject-dependent tasks, while the first session is used for cross-subject tasks.
-III-A2 SEED-IV
+III-D2 SEED-IV
 The SEED-IV[25] dataset comprises 15 subjects (7 males, 8 females), each participating in three sessions with 24 trials per session.
@@
 All sessions’ data are used for subject-dependent tasks, while the first session is used for cross-subject tasks.
-III-A3 DEAP
+III-D3 DEAP
 The DEAP[22] dataset consists of 32 subjects (16 males, 16 females) who participated in one session containing 40 trials.
@@
 All data are used for both subject-dependent and cross-subject tasks.
-III-A4 MAHNOB-HCI
+III-D4 MAHNOB-HCI
 The MAHNOB-HCI[24] dataset includes 30 subjects (13 males, 17 females), though EEG data are completely missing for 2 subjects and partially missing for 3.
@@
 Both complete and partial data are used for subject-dependent and cross-subject tasks.
-III-B Data Preprocessing
-A significant challenge in reproducing extant studies on EER is the lack of a standardized data preprocessing protocol.
-Publicly available EER datasets are not immediately usable and require extensive preprocessing.
-Moreover, this critical information is rarely detailed in the literature, and many open-source packages do not provide the necessary data scripts.
-To address this issue, our benchmark proposes a comprehensive data preprocessing framework specifically designed for EER research.
-The procedure includes: (1) applying bandpass filtering between 0.3 and 50 Hz; (2) eliminating eye movement artifacts using Principal Component Analysis (PCA); (3) extracting DE features across five frequency bands—[0.5, 4], [4, 8], [8, 14], [14, 30], and [30, 50]—followed by processing with a Linear Dynamic System (LDS); and (4) segmenting data using 1-second non-overlapping sliding windows to enhance the dataset.
-III-C Experimental Task
-The field of EER encompasses various experimental tasks, each designed to address distinct challenges with corresponding models.
-The primary experimental tasks are categorized as follows: (1) Subject-dependent: This task evaluates model performance for individual subjects, necessitating both training and testing data from the same subject.
-(2) Cross-subject: This task aims to generalize model performance across different subjects, requiring training and testing data from different subjects.
-(3) Subject-independent: This task assesses model performance across a group of subjects, without needing training and testing data from the same or different subjects.
-(4) Cross-session: This task examines model performance across different sessions of the same subject, requiring training and testing data from different sessions of the same individual.
-The subject-independent task lies between the subject-dependent and cross-subject tasks, with the latter two being more aligned with real-world applications.
-The cross-session task necessitates datasets with multiple sessions, which many datasets lack.
-Therefore, our benchmark focuses on subject-dependent and cross-subject tasks, as these are the most commonly utilized and hold significant practical relevance.
-Additionally, it is important to note that we treat data from different sessions of the same subject as if they originate from distinct subjects.
-Figure 2: The framework and data pipeline of LibEER
-III-D Data Splitting
-Data splitting is a critical process that delineates how a dataset is partitioned into training and test sets, significantly influencing model performance.
-However, prevailing practices in this domain often exhibit a lack of rationality and the absence of a coherent, standardized approach.
-For example, many current data-splitting methodologies fail to incorporate a validation set for model selection and do not consistently adhere to cross-trial configurations, resulting in unreliable experimental outcomes.
-To mitigate these shortcomings, we propose a systematic and rational data-splitting methodology in our benchmark.
-Specifically, for tasks that are subject-dependent, we allocate each individual’s data into training, validation, and test sets in a 0.6:0.2:0.2 ratio, adjusting as necessary for datasets that do not divide evenly.
-In the case of cross-subject tasks, we similarly partition subjects into training, validation, and test sets following the same 0.6:0.2:0.2 ratio.
-Furthermore, our benchmark rigorously adheres to the cross-trial principle, ensuring that samples from the same trial do not appear in both the training and test sets.
-III-E Evaluation Methods
+III-E Evaluation Methods and Metrics
 The evaluation method determines how the results reported in the paper are calculated.
@@
 Therefore, in our benchmark, our evaluation method reports the performance on the test set of the model that achieved the highest F1 score on the validation set across all epochs.
-We use accuracy and F1 score, the most commonly used metrics in existing research, and report both the mean and standard deviation.
+We report both the mean and standard deviation of two metrics, accuracy and F1 score.
+The former is a key metric for classification tasks, while the latter provides a more reasonable assessment when sample labels are imbalanced.
 The methods for calculating accuracy and F1 score are presented in Eq.
 6 and Eq.
-6, respectively.
+7, respectively.
 $\text{ACC}=\frac{TP+TN}{TP+TN+FP+FN}$ (6)
@@
 To facilitate the utilization of the LibEER library in the field of EER, users can easily install via pip, which stremlines the installation process and ensures that all necessary dependencies are managed automatically.
-For further details and access to the source code, please visit the official repository at https://github.com/ButterSen/LibEER.
+For further details and access to the source code, please visit the official repository at https://github.com/XJTU-EEG/LibEER.
 LibEER offers a user-friendly and equitable platform for EER implemented using PyTorch.
@@
 The model training and evaluation module presents a standardized and extensible framework for conducting model training and evaluation.
-Figure 2 illustrates the framework of LibEER.
+Fig.
+3 illustrates the framework of LibEER.
 Below, we provide detailed explanations and usage examples for each module.
@@
 This functionality is illustrated in the usage examples of LibEER for data loading.
-A data loader usage example is presented in Figure 3.
+A data loader usage example is presented in Fig.
+4.
 IV-2 Data split
@@
 To accommodate various splitting methods, the get_split_index() function further divides the data based on the specified formats and labels.
-The splitting methodologies available in LibEER include fixed ratio splitting, which partitions the data into training and testing sets according to a predetermined ratio; early stopping splitting, which divides the data into training, validation, and testing sets while maintaining balanced label distribution; and cross-validation methods, which segment the data into n folds, ensuring that each fold is utilized as the testing set once.
-The following example (see Figure 4) illustrates how to employ LibEER for data splitting.
+The splitting methods available in LibEER include train-test splitting, which partitions the data into training and testing sets according to a predetermined ratio;train-val-test splitting, which divides the data into training, validation, and testing sets while maintaining balanced label distribution;and cross-validation methods, which segment the data into n folds, ensuring that each fold is utilized as the testing set once.
+The following example (see Fig.
+5) illustrates how to employ LibEER for data splitting.
 Figure 6: The example show how to create and train a model
@@
 DBN SEED B, R, DE, 1s dependent 3 : 2 ACC 86.91$\pm$7.62 81.18$\pm$8.13 5.73$\downarrow$
+BiDANN SEED B, R, DE, 9s dependent 3 : 2 ACC 92.38$\pm$7.04 89.06$\pm$9.42 3.32$\downarrow$
+R2G-STNN SEED B, R, DE, 9s dependent 3 : 2 ACC 93.38$\pm$5.96 84.11$\pm$8.47 9.27$\downarrow$
 MS-MDA SEED B, DE, 1s cross 14 : 1 ACC 89.63 93.97 4.34$\uparrow$
@@
 The ultimate outcomes of the model will be reported according to the performance metrics specified by the user.
-The following example in Figure 5 illustrates the process of creating and training a model within LibEER, along with the corresponding evaluation metrics.
-V Experimental Validation
+The following example in Fig.
+6 illustrates the process of creating and training a model within LibEER, along with the corresponding evaluation metrics.
+V Experimental Results
 In this section, we thoroughly validate the importance of LibEER and our benchmark through extensive experimentation.
@@
 Lastly, we conduct ablation studies to address specific issues identified during the experimental phase.
-TABLE V: The Mean Accuracies (and Standard Deviations) using proposed benchmark for Subject-Dependent EEG Emotion Recognition Experiment.
+TABLE V: The Mean Accuracies and F1 scores (and Standard Deviations) using the proposed benchmark for Subject-Dependent EER Experiment.
 The top two methods in each scenario are highlighted using bold and underlined formatting.
@@
 (14.23) (10.79) (14.21) (10.21) (13.15) (8.30) (6.13) (10.92) (12.08) (15.94) (13.52)
-TABLE VI: The Mean Accuracies using proposed benchmark for Cross-Subject EEG Emotion Recognition Experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
+TABLE VI: The Mean Accuracies and F1 scores using the proposed benchmark for Cross-Subject EER Experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
 Method SVM DNN Trans CNN RNN GNN
@@
 Additionally, we detail the specific original experimental setups, clearly illustrating the differences in experimental details among various models.
+All reproduced models have been integrated into our LibEER algorithm library, allowing for direct usage.
 From the results in the Table IV and the replication process, we can draw the following conclusions:
 (1) The reproduced model performances are generally slightly lower than the results reported in the original paper, averaging 2.95% lower, indicating a satisfactory overall replication quality.
-Notably, the reproduction results for RGNN, DBN, GCBNet_BLS, and ACRNN showed more significant declines compared to the original paper.
+Notably, the reproduction results for RGNN, DBN, R2G-STNN, GCBNet_BLS, and ACRNN showed more significant declines compared to the original paper.
 Conversely, some scenarios in MS-MDA, as well as CDCN and HSLT, exhibited improvements in replication results, while the performances of other models were comparable to those in the original study.
@@
 As shown in Table IV, the implementation details of different methods vary significantly, making it difficult to fairly compare their performance and draw reliable conclusions for the advancement of the EER field.
-To this end, we conduct a fair comparison of all methods based on the unified and reasonable benchmark proposed in Section III.
+To this end, we conduct a fair comparison of ten selected methods based on the unified and reasonable benchmark proposed in Section III.
 The results for the two experimental tasks are reported in Tables V and VI, respectively.
@@
 The highest-ranking method among the $n$ methods receives $n$ points, while the lowest-ranking method receives 1 point.
-The total score for each method across all scenarios is then accumulated, and the scoring results for both experimental tasks are presented in Figure 6 (a) and (b).
-From Figure 6 (a), it can be observed that in the subject-dependent task, DGCNN achieved the best performance with a total score of 140.
+The total score for each method across all scenarios is then accumulated, and the scoring results for both experimental tasks are presented in Fig.
+7 (a) and (b).
+From Fig.
+7 (a), it can be observed that in the subject-dependent task, DGCNN achieved the best performance with a total score of 140.
 Following closely are GCBNet and its variant GCBNet_BLS, both scoring 132 points.
@@
 In contrast, EEGNet, Tsception, and ACRNN performed relatively poorly.
-In the cross-subject task (see Figure 6 (b)), the overall performance ranking of each model is similar to that in the subject-dependent scenario.
+In the cross-subject task (see Fig.
+7 (b)), the overall performance ranking of each model is similar to that in the subject-dependent scenario.
 Notably, the DBN model, which has a relatively simple architecture, rises to third place in the cross-subject scenario.
@@
 Several methods performed better on the SEED dataset, which has a more reasonable data collection approach.
-To visually illustrate the performance differences of various models across different datasets, we created radar charts for both experimental tasks, as shown in Figure 7.
-From the results presented in Figure 7, we summarize that models performed better on the three-class SEED dataset compared to the binary classification tasks in DEAP and HCI.
+To visually illustrate the performance differences of various models across different datasets, we created radar charts for both experimental tasks, as shown in Fig.
+8.
+From the results presented in Fig.
+8, we summarize that models performed better on the three-class SEED dataset compared to the binary classification tasks in DEAP and HCI.
 We believe this is due to the higher label quality in SEED, where discrete emotion labels are easier for self-assessment, while continuous VAD labels tend to be more subjective.
@@
 F1 26.63 25.18 23.42 27.70 25.07 18.76
-V-C Discussion of Ablation Study
+V-C Discussion of Experiment Settings
 In this section, we select multiple representative methods from each kind of EER method and conduct ablation studies on three crucial factors, i.e., sample length, feature type, and evaluation methods.
@@
 The ACC and F1 scores for both experimental scenarios are reported.
-V-C1 Sample length
-In the ablation study on sample length, we examined lengths of 1s, 3s, 5s, 7s, and 9s, with the results presented in Figure 8.
-As illustrated in Figure 8, in the subject-dependent scenario, the performance of most methods tends to decline as sample length increases.
+V-C1 Data Splitting Length
+Existing studies have chosen to divide trials into samples of 1 to 10 seconds in length [12, 8, 11].
+To explore the effects and impact of different sample lengths, we examined lengths of 1s, 3s, 5s, 7s, and 9s, with the results presented in Fig.
+9.
+As illustrated in Fig.
+9, in the subject-dependent scenario, the performance of most methods tends to decline as sample length increases.
 We attribute this trend to the sparsity of data in the subject-dependent context; shorter sample lengths yield a greater total number of samples, which enhances the learning capacity of deep models.
@@
 This suggests to researchers that the choice of sample length must balance the trade-off between the quantity of sample augmentation and the quality of the samples.
-V-C2 Feature types
+V-C2 Data preprocessing Feature
 In the ablation study on feature types, we utilized both DE and PSD features, along with their respective variants processed by Locality Differential Sparse (LDS), as inputs.
-The results are presented in Figure 9.
-As illustrated in Figure 9, in the subject-dependent scenario, the DE and PSD features demonstrate comparable performance, while both feature types processed by LDS show significant improvements.
+The results are presented in Fig.
+10.
+As illustrated in Fig.
+10, in the subject-dependent scenario, the DE and PSD features demonstrate comparable performance, while both feature types processed by LDS show significant improvements.
 In the cross-subject scenario, the DE feature exhibited the lowest performance, whereas the other three features displayed similar efficacy.
 This finding underscores the utility of LDS as a valuable feature processing technique, suggesting that the selection of specific features should be adapted flexibly according to the experimental context and the unique characteristics of the model.
-V-C3 Training Methods
+V-C3 Evaluation Methods
 In the ablation study on model training methods, we aimed to explore how different training approaches impact model performance.
@@
 These findings highlight the importance of using a reasonable and appropriate training method for the effective training of EER models.
-VI Limitation
-While LibEER is able to standardize and accelerate research in EER, we are aware of some potential limitations described as follows:
-Incomplete Experimental Settings.
+VI Future Direction
+While LibEER is able to standardize and accelerate research in EER, we will further improve and enhance the following two aspects in our future work.
+Model Reproduction.
+In replicating representative models in the EER field, we faced several challenges despite closely following the papers.
+Some methods lacked detailed descriptions of model or training parameters[5, 7].
+In other cases [12, 13, 11], the available code did not fully align with the reported paper details.
+We also found that in DGCNN[12] minor adjustments during replication improved performance, suggesting the paper’s content may not have been updated.
+When the above issues arose, we used our experience to replicate the models as accurately as possible, adhering to the papers.
+The replication results (Table IV) show discrepancies compared to the original reports.
+In the future, we aim to collaborate with original authors and the broader EER community to improve and expand this algorithm library.
+Experimental Settings.
 LibEER includes several key experimental setups to facilitate research, including four widely recognized public datasets, common data preprocessing techniques, two main tasks, common data partitioning methods, and reasonable performance evaluation methods.
@@
 Future versions of LibEER will aim to incorporate more diverse experimental setups from the EER field.
-Inadequate Model Reproduction.
-In replicating representative models in the EER field, we faced several challenges despite closely following the papers.
-Some methods lacked detailed descriptions of model or training parameters[5, 7].
-In other cases [12, 13, 11], the available code did not fully align with the reported paper details.
-We also found that in DGCNN[12] minor adjustments during replication improved performance, suggesting the paper’s content may not have been updated.
-When the above issues arose, we used our experience to replicate the models as accurately as possible, adhering to the papers.
-The replication results (Table IV) show discrepancies compared to the original reports.
-In the future, we aim to collaborate with original authors and the broader EER community to improve and expand this algorithm library.
 VII Conclusion

### v3

@@
 LibEER carefully selects popular and powerful baselines, harmonizes key implementation details across methods, and provides a standardized codebase in PyTorch.
-By offering a consistent evaluation framework with standardized experimental settings, LibEER enables unbiased assessments of over ten representative deep learning models for EER across the four most widely used datasets.
+By offering a consistent evaluation framework with standardized experimental settings, LibEER enables unbiased assessments of seventeen representative deep learning models for EER across the six most widely used datasets.
 Additionally, we conduct a thorough, reproducible comparison of model performance and efficiency, providing valuable insights to guide researchers in the selection and design of EER models.
@@
 I Introduction
-Emotion represents a spectrum of subjective cognitive experiences that forms the core framework for interpersonal relationships and serves as a foundation for evaluating human behaviors [1].
-Due to its non-invasive nature, high temporal resolution, portability, and cost-effectiveness, electroencephalography (EEG) has become a prominent physiological signal in emotion analysis [30, 31].
-As a result, EEG-based emotion recognition (EER) has attracted significant research interest in recent years, with applications spanning diverse fields such as healthcare [32, 41], advertising [33], and education [34].
-Traditional approaches to EER primarily rely on hand-crafted features, such as power spectral density (PSD)[40] and differential entropy (DE)[16].
-Recently, with advancements in automatic feature engineering, deep learning models have emerged as a promising research area in EER and achieved dominant performance [17, 19, 2].
-These models generally target to extract temporal features within each EEG channel, spatial features across different EEG channels, and spectral features in the frequency domain.
+Emotion represents a spectrum of subjective cognitive experiences that underpins interpersonal relationships and human behavior assessment [1].
+Electroencephalography (EEG), owing to its non-invasive nature, high temporal resolution, portability, and cost-effectiveness, has become a widely used physiological signal in emotion analysis [37, 38].
+Consequently, EEG-based emotion recognition (EER) has garnered increasing attention across domains such as healthcare [39, 48], advertising [40], and education [41].
+Early EER methods focused on hand-crafted features like power spectral density (PSD)[47] and differential entropy (DE)[21], while recent advances in deep learning have enabled automatic extraction of temporal, spatial, and spectral features, leading to state-of-the-art performance [22, 24, 2].
 Despite significant advancements in deep learning model design, the practical implementation of these approaches is often overlooked, making it difficult for newcomers and researchers to keep up with emerging studies.
@@
 Therefore, there is a pressing need for an open-source library based on a standardized platform that provides comprehensive implementation details while achieving performance comparable to the original study.
-  • I3 (Lack of in-depth analysis of SOTA EER research): While deep learning has dominated EER research, many survey papers  have been published summarizing progress from various perspectives.
+  • I3 (Lack of in-depth analysis of SOTA EER research): While deep learning has dominated EER research, many survey papers [2, 4] have been published summarizing progress from various perspectives.
 However, these surveys typically report findings without hands-on experience in reproducing results, which may lead to misdirected future research.
@@
 Therefore, an in-depth analysis of SOTA EER research based on reproduction results is essential for uncovering the true challenges and advancing the field.
-In response to these challenges, we propose LibEER (Library for EEG-based Emotion Recognition) and conduct an analysis of SOTA EER research based on reproduction results with LibEER.
-To establish a comprehensive benchmark in EER (addressing issue I1), we standardize datasets, evaluation metrics, and experimental settings.
-Specifically, we review recent SOTA EER studies to identify the most commonly used datasets, SEED [23], SEED-IV [25], DEAP [22], and MAHNOB-HCI [24], and evaluate the corresponding metrics.
-Our benchmark includes both subject-dependent and subject-independent settings, along with a variety of preprocessing methods.
-In contrast to the simplified train/test split commonly used in the literature, we adopt a more rigorous train/validation/test split to ensure a more reliable performance evaluation.
-To address issue I2, we develop the LibEER by systematically selecting and implementing ten representative deep learning-based EER methods, all within a unified training pipeline in PyTorch.
-LibEER offers a wide range of EEG signal preprocessing functions, evaluation protocols, and experimental scenarios.
-To enhance its flexibility, we provide user-friendly interfaces that allow researchers to easily customize experimental configurations.
-Users can set hyperparameters, choose preprocessing methods, and modify backbone models to suit their specific needs.
-Based on the reproduction results from LibEER, we conduct a series of comprehensive comparison experiments and identify four key observations (addressing issue I3).
-Through these experiments, we highlight both the real limitations and the progress made in current studies.
-It is worth noting that recent research [3] has made important contributions by providing a toolbox for EER.
-However, our work differs significantly and builds upon these contributions in several key areas.
-First, we systematically analyze the challenges associated with making fair comparisons in EER tasks, a discussion notably absent in [3].
-Second, while the existing benchmark [3] includes a variety of methods, many of them are not specifically tailored for EER tasks.
-In contrast, we have selected the ten most representative deep learning-based EER methods, based on their impact in this field, making our benchmark better reflect the true SOTA.
-Finally, the existing benchmark [3] predominantly focuses on research that has already made its code publicly available.
-However, many influential EER studies do not provide open-source code, which has led to the exclusion of some significant models in their work.
-In response, we have carefully reproduced all representative studies, including those without publicly available code and rigorously validated the results to ensure their accuracy.
-We believe that our benchmark and library offer fresh perspectives that will significantly advance the field.
+To address these issues, we propose LibEER (Library for EEG-based Emotion Recognition), an open-source platform that enables reproduction and analysis of SOTA EER studies.
+To establish a benchmark (addressing I1), we standardize datasets, evaluation metrics, and experimental settings.
+Based on a review of recent studies, we include six widely used datasets, SEED [28], SEED-IV [30], DEAP [27], MAHNOB-HCI [29], SEED-V [31], and MPED [32], and support both subject-dependent and cross-subject configurations, using a rigorous train/validation/test split.
+To address I2, we implement seventeen representative deep learning-based EER methods within a unified PyTorch training pipeline.
+LibEER provides diverse EEG preprocessing tools, evaluation protocols, and flexible experimental configurations via user-friendly interfaces.
+Finally, to address I3, we conduct extensive reproduction-based comparisons using LibEER and derive four key observations that reveal the strengths and limitations of existing models.
+While recent efforts like [3] have introduced EER toolkits, our work differs by providing more in-depth benchmarking analysis, explicitly focusing on EER-specific models, and systematically reproducing several influential methods that lack open-source implementations.
+These contributions together offer a more comprehensive and practical foundation for advancing EER research.
 In summary, our main contributions are as follows:
   • We have pioneered the establishment of the first benchmark specifically designed to enable fair comparisons in EEG-based emotion recognition (EER).
-This benchmark not only reveals the true progress of the EER task but also provides a convenient reference for conducting comparative experiments.
+This benchmark not only offers an objective reflection of the actual progress in the EER field but also serves as a practical reference for conducting standardized comparative experiments.
   • We have introduced LibEER as an open-source project on GitHub, offering flexible customization of experimental schemes through user-friendly interfaces.
-LibEER supports fair and comprehensive evaluation of over ten popular models across four datasets.
+LibEER supports fair and comprehensive evaluation of seventeen popular models across six datasets.
 Serving both as a toolbox and a platform, it assists practitioners in efficiently utilizing and reproducing EER methods.
@@
 The paper is organized as follows.
-Section II provides the background of the EER task.
+Section II provides the background and formal definition of the EER task.
 Section III presents the design of the benchmark.
@@
 1.
-Each data collection begins with a hint of start shown to the participant.
+Each data collection begins with a start hint shown to the participant.
 Afterwards, various stimuli, such as images, audio, or video, are presented to evoke specific emotions in the participant, such as happiness, sadness, or fear.
@@
 Two consecutive segments within a trial can be overlapping or non-overlapping.
-Each segment is called a sample, which serves as the fundamental unit for EER.
+Each segment is called a sample, serving as the fundamental unit for EER.
 Notably, within a single trial, all samples should be designated for either training or test.
-A trial-dependent setting, where some samples within a trial are used for training and others for testing, is impractical for real-world applications and should be considered an invalid setup.
+Splitting samples from the same trial across training and test sets is impractical for real-world applications and considered invalid.
 II-B Problem Formulation
@@
 Experimental Task Subject-dependent / Cross-subject
-Datasets SEED / SEED-IV / DEAP / MAHNOB-HCI
+Datasets SEED / SEED-IV / DEAP / MAHNOB-HCI / SEED-V / MPED
 Evaluation Metrics Accuracy / F1-score
-Figure 2: The distinct utilization of EEG signals by different methods
+Figure 2: The distinct utilization of EEG signals by different methods.
 TABLE II: Detailed information of selected baseline methods.
-Model Baseline frequency Citation count Publication Publish year Type
+Model Base. freq. Citation count Publication Publish year Type
+EEGNet[11] 7 2979 JNE 2018 CNN
+CDCN[12] 2 142 TCDS 2020 CNN
 Tsception[13] 2 136 TAFFC 2022 CNN
-CDCN[9] 2 142 TCDS 2020 CNN
-EEGNet[8] 7 2979 JNE 2018 CNN
+ACRNN[14] 3 338 TAFFC 2020 RNN
+BiDANN[43] 4 250 TAFFC 2021 RNN
 R2G-STNN[17] 8 235 TAFFC 2022 RNN
-BiDANN[36] 4 250 TAFFC 2021 RNN
-ACRNN[11] 3 338 TAFFC 2020 RNN
-MS-mda[6] 6 95 Front. Neurosci. 2021 DNN
+DGCNN[15] 18 1105 TAFFC 2018 GNN
+GCBNet[18] 3 302 TAFFC 2019 GNN
+GCBNet_BLS[18] 3 302 TAFFC 2019 GNN
+RGNN[16] 11 545 TAFFC 2020 GNN
+NSAL-DGAT [20] - 1 TAFFC 2025 GNN
 DBN[6] 7 1884 TAMD 2015 DNN
-RGNN[13] 11 545 TAFFC 2020 GNN
-GCBNet[15] 3 302 TAFFC 2019 GNN
-DGCNN[12] 18 1105 TAFFC 2018 GNN
-HSLT[7] 2 79 IEEE SENS J 2022 Transformer
+DAN [8] - 199 ICONIP 2018 DNN
+DANN [8] - 199 ICONIP 2018 DNN
+MS-MDA[7] 6 95 Front. Neuros. 2021 DNN
+PR-PL [9] - 55 TAFFC 2024 DNN
+HSLT[10] 2 79 IEEE SENS J 2022 Trans.
 III Benchmark Building
@@
   • After that, we ranked these 205 methods based on their frequency as comparison methods in the 60 papers and selected the top 20 for replication.
-  • These 20 methods, based on model architecture, are divided into five categories: CNN, RNN, GNN, DNN, and Transformer.
-Finally, we selected models with replication results within a 10% error margin compared to the original reported results, while ensuring that each category includes at least one representative model.
+  • These 20 methods, following the common practice in previous studies of categorizing deep learning models by their main architecture [66, 67], are divided into five categories: CNN, RNN, GNN, DNN, and Transformer.
+Finally, we selected models with replication results within a 10% error margin compared to the original reported results, ensuring that each category includes at least one representative model.
 Based on these steps, we ultimately identified the twelve most representative works [11, 12, 13, 14, 15, 16, 18, 6, 7, 10, 43, 17] to include in LibEER.
-TABLE II presents detailed information on these twelve methods.
+In addition, we include two classic transfer learning methods (DAN and DANN [8]) that have demonstrated superior performance in EER, as well as two recent state-of-the-art approaches (NSAL-DGAT [20] and PR-PL [9]).
+These methods are also categorized into the corresponding architecture-based groups according to their backbone model structures.
+TABLE II presents detailed information on these seventeen methods.
 In future work, we will continue to maintain and expand this collection by incorporating more recent and diverse methods.
@@
 LibEER includes three representative CNN-based methods: EEGNet [11], CDCN [12], and TSception [13].
-Although not initially proposed for EER, EEGNet is included as a representative 2D CNN-based model due to its significant influence, with over 3000 citations.
-Both CDCN and TSception employ 1D CNNs for spatial and temporal feature extraction.
-TSception, a more advanced approach, incorporates multi-scale 1D CNNs for temporal features and employs hemisphere kernels for spatial features.
 III-A2 RNN-based Methods
@@
 Our benchmark includes three RNN-based methods: ACRNN [14], BiDANN [43], and R2G-STNN [17].
-Both ACRNN and BiDANN use LSTM for temporal feature extraction, while R2G-STNN applies bidirectional LSTM to capture spatial features.
-An attention mechanism is often integrated with LSTM to enhance feature extraction; for instance, ACRNN uses channel-wise attention for spatial features and extended self-attention with LSTM for temporal feature extraction.
 III-A3 GNN-based Methods
@@
 Since EEG data does not inherently contain information about the edges or the adjacency matrix, using graph neural networks to process EEG data often involves leveraging prior knowledge about the relationships between channels or using learnable adjacency matrices as the graph’s adjacency matrix.
-We introduced three GNN-based methods.
-Dynamical Graph CNN (DGCNN)  [12] models multichannel EEG features using a dynamic adjacency matrix that adapts during training.
-It employs Chebyshev graph convolution for efficient filtering calculations.
-Regularized GNN (RGNN)  [13] utilizes a neuroscience-inspired adjacency matrix and introduces simple graph convolution networks (SGC).
-RGNN enhances robustness through node-wise domain adversarial training and emotion-aware distribution learning to address cross-subject variations and noisy labels.
-Graph Convolutional Broad Network (GCBNet)  [15] combines graph convolutional layers with traditional convolutional layers for effective feature extraction.
-GCBNet uses Chebyshev graph convolution followed by stacked convolutional layers to abstract high-level features.
-The GCBNet+BLS variant incorporates the Broad Learning System to enhance feature extraction and classification further.
+Our benchmark includes five GNN-based methods: DGCNN [15], GCBNet [18], GCBNet_BLS [18], RGNN [16], and NSAL-DGAT [20].
 III-A4 DNN-based Methods
-Deep Neural Network (DNN) is a feedforward artificial network consisting of fully connected neurons with a nonlinear activation function, including the input layer, multiple hidden layers, and the output layer.
-In the field of EER, DNN processes EEG data by transforming and mapping it into the feature space, enabling the extraction and utilization of complex patterns within the data, illustrated in Fig.
+Deep Neural Network (DNN), in this paper, specifically refers to the feedforward fully connected neural network, also known as Fully Connected Neural Network (FNN) or Multi-Layer Perceptron (MLP).
+It consists of an input layer, multiple hidden layers composed of fully connected neurons with nonlinear activation functions, and an output layer.
+Unlike other deep learning architectures such as CNN, RNN, or GNN, DNN focuses on dense layer transformations without convolutional, recurrent, or graph-based structures.
+In the field of EER, DNN processes EEG data by performing feature transformation and mapping into a high-dimensional feature space, enabling the extraction and utilization of complex patterns within the data, as illustrated in Fig.
 2 (d).
-In this paper, we have implemented two popular DNN-based methods.
-Deep Belief Networks (DBN) [5] are probabilistic generative models representing input data distributions through hidden variables.
-They consist of stacked restricted Boltzmann machines (RBMs), with the output of one RBM serving as input for the next, and use an efficient greedy layer-wise algorithm for pre-training.
-Multi-Source Marginal Distribution Adaptation (MS-MDA) [6] utilizes MLPs for feature extraction and classification while employing domain adaptation to reduce variance across subjects and sessions.
-MS-MDA maintains independent branches rather than merging source domains, aligning each with the target domain individually to preserve marginal distributions and enhance adaptation performance.
-III-A5 Transform-based Methods
+Our benchmark includes five DNN-based methods: DBN [6], DAN [8], DANN [8], MS-MDA [7], and PR-PL [9].
+III-A5 Transformer-based Methods
 Transformer uses a self-attention mechanism to capture dependencies between elements in a sequence, effectively handling long-range dependencies.
@@
 These patches are then fed into the transformer model for further analysis.
-The Hierarchical Spatial Learning Transformer (HSLT) [7] is a hierarchical transformer that extracts discriminative features by capturing EEG spatial dependencies from electrode to brain-region level.
-It inputs electrode-extracted patches into the first layer of the transformer encoder to aggregate information within the same brain region.
-The output then serves as input for the second layer, which captures spatial dependencies between different brain regions and highlights those with greater contributions.
+Our benchmark includes one Transformer-based methods: HSLT [10].
 III-B Data preprocessing and Splitting
@@
 Specifically, for tasks that are subject-dependent, we allocate each individual’s data into training, validation, and test sets in a 0.6:0.2:0.2 ratio, adjusting as necessary for datasets that do not divide evenly.
+The SEED-V dataset contains relatively fewer trials and more classes, so the ratio was adjusted to 1:1:1.
 In the case of cross-subject tasks, we similarly partition subjects into training, validation, and test sets following the same 0.6:0.2:0.2 ratio.
@@
 The field of EER encompasses various experimental tasks, each designed to address distinct challenges with corresponding models.
-The primary experimental tasks are categorized as follows: (1) Subject-dependent: This task evaluates model performance for individual subjects, necessitating both training and testing data from the same subject.
+The primary experimental tasks are categorized as follows: (1) Subject-dependent: This task evaluates model performance for individual subjects, requiring both training and testing data from the same subject.
 (2) Cross-subject: This task aims to generalize model performance across different subjects, requiring training and testing data from different subjects.
-(3) Subject-independent: This task assesses model performance across a group of subjects, without needing training and testing data from the same or different subjects.
+(3) Subject-independent: This task assesses model performance without constraining the subject-level split, allowing training and testing data to be randomly divided from mixed samples of multiple subjects.
 (4) Cross-session: This task examines model performance across different sessions of the same subject, requiring training and testing data from different sessions of the same individual.
-The subject-independent task lies between the subject-dependent and cross-subject tasks, with the latter two being more aligned with real-world applications.
-The cross-session task necessitates datasets with multiple sessions, which many datasets lack.
-Therefore, our benchmark focuses on subject-dependent and cross-subject tasks, as these are the most commonly utilized and hold significant practical relevance.
+Figure 3 provides an intuitive illustration of the differences among the four tasks.
+Figure 3: Four different experimental tasks in EER.
+In previous studies, the definitions of subject-independent and cross-subject tasks have often overlapped, as subject-independent is typically implemented using leave-one-subject-out cross-validation (LOSO-CV) [15, 16].
+To avoid ambiguity, we explicitly adopt cross-subject as one of the core evaluation settings in our benchmark.
+Cross-subject strictly requires training and testing on entirely different subjects, emphasizing the model’s generalization ability.
+This task setup is more aligned with real-world applications where models are expected to perform reliably on new, unseen users in large-scale deployment scenarios.
+In addition, subject-dependent tasks are included to evaluate a model’s performance under personalized settings, where training and testing are performed on the same subject.
+This setup is important for scenarios focusing on optimizing accuracy for specific individuals, such as personalized healthcare or user-specific emotion monitoring.
+Therefore, our benchmark concentrates on cross-subject and subject-dependent tasks as the most widely adopted and practically meaningful evaluation setups.
 Additionally, it is important to note that we treat data from different sessions of the same subject as if they originate from distinct subjects.
@@
 MAHNOB-HCI[29] 1 30 20 32 movie clip 34.9-117s 512 raw data valence/arousal
-Figure 3: The framework and data pipeline of LibEER
+SEED-V [31] 3 16 15 62 movie clip 120s-240s 1000 raw data happy/sad/disgust/netural/fear
+MPED [32] 1 23 28 62 movie clip 150s-300s 1000 STFT joy/funny/anger/fear/disgust/sad/neutrality
 III-D Datasets
 Research in the EER field exhibits variability in dataset selection and often lacks clarity regarding the specific data segments utilized.
-This study evaluates the performance of popular models using four datasets, which are selected by comprehensively considering factors such as their frequency of use, data scale, and data quality.
+This study evaluates the performance of popular models using six datasets, which are selected by comprehensively considering factors such as their frequency of use, data scale, and data quality.
 They will be introduced along with the relevant data segments, as summarized in Table III.
@@
 EEG data were recorded using a 32-channel cap at 512 Hz, and the dataset includes both raw and downsampled (128 Hz) versions.
-All data are used for both subject-dependent and cross-subject tasks.
+All samples are used for both subject-dependent and cross-subject tasks.
+In the subsequent experiments, we used the DEAP dataset with the valence labels, the DEAP dataset with the arousal labels, and the DEAP dataset with both labels, denoted as DEAP-V, DEAP-A, and DEAP-VA, respectively.
 III-D4 MAHNOB-HCI
@@
 EEG data were recorded using a 32-channel system at 512 Hz, and the dataset includes both raw and downsampled (128 Hz) versions.
-Both complete and partial data are used for subject-dependent and cross-subject tasks.
+Both complete and partial samples are used for subject-dependent and cross-subject tasks.
+In the subsequent experiments, we used the MAHNOB-HCI dataset with the valence labels, the MAHNOB-HCI dataset with the arousal labels, and the MAHNOB-HCI dataset with both labels, denoted as HCI-V, HCI-A, and HCI-VA, respectively.
+III-D5 SEED-V
+The SEED-V[31] dataset contains 16 subjects (6 males, 10 females) who each participated in three experimental sessions.
+Each session included 15 emotion-inducing movie clips (3 clips per emotion category: happy, sad, neutral, fear, and disgust).
+EEG signals were recorded using a 62-channel cap following the 10-20 system, accompanied by simultaneous eye movement recordings.
+The dataset provides differential entropy (DE) features extracted from five frequency bands (delta, theta, alpha, beta, and gamma) for EEG signals.
+Data from all three sessions are used for subject-dependent and cross-subject tasks.
+III-D6 MPED
+The MPED[32] dataset is a comprehensive multimodal resource for discrete emotion recognition, comprising physiological recordings from 23 Chinese participants (10 males, 13 females).
+During emotion elicitation experiments using 28 standardized video stimuli representing seven discrete emotions (joy, funny, anger, sadness, disgust, fear, neutrality), 62-channel EEG signals were acquired via an ESI NeuroScan System at 1000 Hz sampling rate following the international 10-20 system.
+All data are used for both subject-dependent and cross-subject tasks.
 III-E Evaluation Methods and Metrics
@@
 where TP (True Positives), TN (True Negatives), FP (False Positives), and FN (False Negatives) are the metrics used to calculate accuracy and F1 score.
+Besides, the random seed is fixed at 2024 to ensure the reproducibility of the results.
 IV LibEER Toolkit Framework
@@
 For further details and access to the source code, please visit the official repository at https://github.com/XJTU-EEG/LibEER.
+Figure 4: The framework and data pipeline of LibEER.
 LibEER offers a user-friendly and equitable platform for EER implemented using PyTorch.
@@
 4 illustrates the framework of LibEER.
-Below, we provide detailed explanations and usage examples for each module.
-Figure 4: The example shows how to LibEER to load data
-Figure 5: The example how to use LibEER to slit data
+Below, we provide detailed explanations for each module.
 IV-1 Data loader
@@
 Ultimately, the processed data is integrated into the format of (session, subject, trial, sample).
-This functionality is illustrated in the usage examples of LibEER for data loading.
-A data loader usage example is presented in Fig.
-4.
+An example demonstrating how to use the data loader is provided in our GitHub project.
 IV-2 Data split
@@
 The splitting methods available in LibEER include train-test splitting, which partitions the data into training and testing sets according to a predetermined ratio;train-val-test splitting, which divides the data into training, validation, and testing sets while maintaining balanced label distribution;and cross-validation methods, which segment the data into n folds, ensuring that each fold is utilized as the testing set once.
-The following example (see Fig.
-5) illustrates how to employ LibEER for data splitting.
-Figure 6: The example show how to create and train a model
+An example illustrating how to perform data splitting with LibEER is provided in our GitHub project.
 TABLE IV: Experimental settings and comparison of replication results for reproduced models in LibEER.
@@
 Dataset Preprocessing Task Splitting  (Train : Test) Evaluation Reported (%) Ours (%) Gap (%)
-DGCNN SEED B, R, DE, 1s dependent 3 : 2 ACC 90.40$\pm$8.49 89.48$\pm$8.49 0.92$\downarrow$
-RGNN SEED B, R, DE, 1s dependent 3 : 2 ACC 94.24$\pm$5.95 84.66$\pm$10.74 9.58$\downarrow$
 EEGNet SEED B, R, DE, 1s dependent 3 : 2 ACC —— 68.15$\pm$12.32 ——
-DBN SEED B, R, DE, 1s dependent 3 : 2 ACC 86.91$\pm$7.62 81.18$\pm$8.13 5.73$\downarrow$
-BiDANN SEED B, R, DE, 9s dependent 3 : 2 ACC 92.38$\pm$7.04 89.06$\pm$9.42 3.32$\downarrow$
-R2G-STNN SEED B, R, DE, 9s dependent 3 : 2 ACC 93.38$\pm$5.96 84.11$\pm$8.47 9.27$\downarrow$
-MS-MDA SEED B, DE, 1s cross 14 : 1 ACC 89.63 93.97 4.34$\uparrow$
-GCBNet SEED B, R, DE, 1s dependent 3 : 2 ACC 92.30$\pm$7.40 89.04$\pm$8.03 3.26$\downarrow$
-GCBNet_BLS SEED B, R, DE, 1s dependent 3 : 2 ACC 94.24$\pm$6.70 88.80$\pm$9.54 5.44$\downarrow$
 CDCN SEED B, R, DE, 1s dependent 3 : 2 ACC 90.63 85.10$\pm$8.80 5.53$\downarrow$
@@
 DEAP-A B, R, 1s dependent 9 : 1 ACC 92.92 91.99$\pm$12.20 0.93$\downarrow$
+TSception DEAP-V B, R, 4s dependent 9 : 1 ACC 61.57$\pm$11.04 61.89$\pm$12.16 0.32$\uparrow$
+DEAP-A B, R, 4s dependent 9 : 1 ACC 59.14$\pm$7.60 60.02$\pm$8.84 0.88$\uparrow$
 ACRNN DEAP-V B, R, DE, 3s dependent 9 : 1 ACC 93.72 86.03$\pm$9.20 7.69$\downarrow$
 DEAP-A B, R, DE, 3s dependent 9 : 1 ACC 93.38 88.31$\pm$7.77 5.07$\downarrow$
-HSLT DEAP-V B, R, 6s cross 14 : 1 ACC, F1 66.51 69.18 2.67$\uparrow$
-DEAP-A B, R, 6s cross 14 : 1 ACC, F1 65.75 68.81 3.06$\uparrow$
-DEAP-VA B, R, 6s cross 14 : 1 ACC, F1 56.93 49.57 7.36$\downarrow$
+BiDANN SEED B, R, DE, 9s dependent 3 : 2 ACC 92.38$\pm$7.04 89.06$\pm$9.42 3.32$\downarrow$
+SEED B, R, DE, 9s cross 14 : 1 ACC 83.28$\pm$9.60 80.36$\pm$8.25 2.92$\downarrow$
+R2G-STNN SEED B, R, DE, 9s dependent 3 : 2 ACC 93.38$\pm$5.96 88.71$\pm$8.85 4.67$\downarrow$
+SEED B, R, DE, 9s cross 14 : 1 ACC 84.16$\pm$7.63 78.35$\pm$9.46 5.81$\downarrow$
+DGCNN SEED B, R, DE, 1s dependent 3 : 2 ACC 90.40$\pm$8.49 89.48$\pm$8.49 0.92$\downarrow$
+GCBNet SEED B, R, DE, 1s dependent 3 : 2 ACC 92.30$\pm$7.40 89.04$\pm$8.03 3.26$\downarrow$
+GCBNet_BLS SEED B, R, DE, 1s dependent 3 : 2 ACC 94.24$\pm$6.70 88.80$\pm$9.54 5.44$\downarrow$
+RGNN SEED B, R, DE, 1s dependent 3 : 2 ACC 94.24$\pm$5.95 84.66$\pm$10.74 9.58$\downarrow$
+NSAL-DGAT SEED B, R, DE, 1s dependent 3 : 2 ACC 96.53$\pm$9.10 92.27$\pm$10.09 4.26$\downarrow$
+SEED B, R, DE, 1s cross 14 : 1 ACC 95.69$\pm$5.24 90.19$\pm$7.32 5.50$\downarrow$
+DBN SEED B, R, DE, 1s dependent 3 : 2 ACC 86.91$\pm$7.62 81.18$\pm$8.13 5.73$\downarrow$
+DAN SEED B, R, DE, 1s cross 14 : 1 ACC 83.81$\pm$8.56 76.29$\pm$6.60 7.52$\downarrow$
+DANN SEED B, R, DE, 1s cross 14 : 1 ACC —— 77.40$\pm$5.86 ——
+MS-MDA SEED B, DE, 1s cross 14 : 1 ACC 89.63$\pm$6.79 89.65$\pm$7.24 0.02$\uparrow$
+PR-PL SEED B, R, DE, 1s dependent 3 : 2 ACC 94.84$\pm$9.16 93.2$\pm$6.66 1.64$\downarrow$
+SEED B, R, DE, 1s cross 14 : 1 ACC 94.84$\pm$9.16 89.19$\pm$5.34 5.65$\downarrow$
+HSLT DEAP-V B, R, PSD, 6s cross 14 : 1 ACC, F1 66.51 69.18 2.67$\uparrow$
+DEAP-A B, R, PSD, 6s cross 31 : 1 ACC, F1 65.75 68.81 3.06$\uparrow$
+DEAP-VA B, R, PSD, 6s cross 31 : 1 ACC, F1 56.93 49.57 7.36$\downarrow$
 B: Bandpass filtering, R: Removing eye movement, DE: Extracting DE features, Ns: Segmenting data into N seconds.
@@
 In scenarios where a validation set is not available, LibEER identifies the optimal performance on the testing set during training as the final outcome for that training iteration.
-Conversely, when a validation set is utilized, LibEER selects the network weights that yield the best performance on the validation set for evaluation against the testing set, thereby determining the final result of the training round.
+Conversely, when a validation set is utilized, LibEER selects the network weights that yield the best performance on the validation set for evaluation against the testing set, thereby determining the final result.
 The ultimate outcomes of the model will be reported according to the performance metrics specified by the user.
-The following example in Fig.
-6 illustrates the process of creating and training a model within LibEER, along with the corresponding evaluation metrics.
+An example of creating, training, and evaluating a model within LibEER is provided in our GitHub project.
 V Experimental Results
@@
 Subsequently, we employ LibEER to compare the performance of all models based on the established benchmark.
-Lastly, we conduct ablation studies to address specific issues identified during the experimental phase.
+Lastly, we investigate the impact of several key experimental settings on model performance through comparative experiments.
 TABLE V: The mean accuracies and F1 scores (and standard deviations) using the proposed benchmark for subject-dependent EER experiment.
 The top two methods in each scenario are highlighted using bold and underlined formatting.
-Method SVM DNN Trans CNN RNN GNN
-DBN HSLT EEGNet CDCN TSception ACRNN DGCNN RGNN GCBNet GCBNet_BLS
-SEED ACC 75.08 71.88 64.83 58.81 68.23 64.01 49.71 82.55 76.55 80.56 76.64
-(19.73) (19.02) (20.47) (16.22) (20.35) (16.44) (13.15) (15.61) (16.92) (16.98) (17.44)
-F1 70.82 67.39 58.82 54.41 63.76 60.53 45.78 79.89 72.52 77.29 72.52
-(23.51) (22.81) (23.36) (17.59) (24.49) (18.51) (14.18) (18.93) (20.08) (20.92) (21.20)
-SEED-IV ACC 47.80 45.56 40.28 29.89 52.26 36.06 29.01 52.39 45.40 53.28 53.51
-(23.03) (21.19) (23.80) (13.53) (21.97) (15.12) (7.10) (24.32) (22.90) (21.05) (22.45)
-F1 40.17 37.61 30.92 26.59 45.26 32.77 19.80 45.94 38.24 46.26 46.91
-(21.68) (20.68) (24.47) (13.58) (23.00) (15.08) (5.42) (24.17) (23.09) (22.27) (22.46)
-HCI-V ACC 64.83 62.03 64.00 61.15 60.48 61.12 60.51 67.83 64.86 66.84 69.60
-(25.95) (24.90) (11.40) (16.76) (21.90) (15.52) (16.89) (22.40) (17.36) (21.42) (22.09)
-F1 55.99 52.84 55.77 50.35 51.73 50.51 49.39 54.78 50.41 54.61 57.78
-(27.80) (27.04) (13.12) (17.28) (23.11) (16.69) (15.70) (26.69) (20.34) (25.08) (27.18)
-HCI-A ACC 63.61 68.51 67.74 67.42 71.82 68.26 66.26 67.29 70.96 64.89 69.60
-(23.44) (21.72) (17.22) (21.71) (21.72) (23.10) (22.69) (27.73) (19.79) (27.12) (22.09)
-F1 50.99 57.18 58.42 54.50 62.89 56.29 55.17 58.04 57.66 57.43 57.78
-(24.89) (26.63) (19.50) (20.05) (24.99) (23.79) (23.20) (31.41) (25.39) (29.50) (27.18)
-HCI-VA ACC 46.29 44.38 46.99 38.32 52.00 40.00 41.00 53.06 49.46 49.15 49.24
-(28.42) (26.78) (20.76) (19.51) (26.05) (20.60) (21.58) (24.44) (23.20) (26.16) (27.97)
-F1 33.91 31.96 34.76 24.56 38.04 27.19 27.10 39.75 35.97 36.49 36.68
-(27.65) (27.87) (19.69) (13.71) (26.88) (13.65) (15.13) (26.01) (23.84) (27.58) (27.25)
-DEAP-V ACC 54.17 56.08 56.20 51.50 57.71 51.52 53.52 56.07 55.90 56.49 57.02
-(18.67) (17.38) (18.14) (11.57) (14.72) (9.54) (9.29) (17.15) (16.24) (18.17) (15.07)
-F1 49.73 48.61 48.21 47.85 53.41 47.33 48.31 49.08 47.25 50.36 51.40
-(18.92) (19.33) (18.73) (11.70) (15.50) (9.58) (7.77) (17.50) (17.55) (19.57) (17.25)
-DEAP-A ACC 63.49 64.60 59.74 61.30 63.37 57.49 61.83 62.68 66.09 65.95 61.07
-(16.72) (19.42) (18.82) (15.88) (14.18) (11.86) (14.32) (19.66) (13.91) (17.61) (16.56)
-F1 53.31 52.61 50.10 53.26 53.94 50.75 49.68 53.94 49.27 55.34 50.43
-(14.39) (19.85) (18.08) (13.05) (13.76) (11.30) (9.20) (20.10) (12.86) (17.86) (16.36)
-DEAP-VA ACC 37.32 39.50 43.34 39.41 38.08 35.68 38.20 41.86 44.53 38.80 37.51
-(17.24) (13.99) (14.49) (11.53) (15.52) (12.08) (11.34) (11.57) (14.35) (16.23) (15.29)
-F1 25.55 24.88 23.47 29.19 28.90 26.75 21.05 29.12 25.88 27.89 27.00
-(14.23) (10.79) (14.21) (10.21) (13.15) (8.30) (6.13) (10.92) (12.08) (15.94) (13.52)
-TABLE VI: The Mean Accuracies and F1 scores using the proposed benchmark for Cross-Subject EER Experiment. The top two methods in each scenario are highlighted using bold and underlined formatting.
-Method SVM DNN Trans CNN RNN GNN
-DBN MS-MDA HSLT EEGNet CDCN TSception ACRNN DGCNN RGNN GCBNet GCBNet_BLS
-SEED ACC 37.07 36.16 64.00 56.00 38.19 57.72 45.60 45.39 60.87 57.20 56.32 56.32
-F1 33.45 22.67 57.35 55.75 31.83 58.66 43.54 42.37 57.22 51.37 55.12 51.43
-SEED-IV ACC 28.98 36.82 56.07 30.33 28.19 31.03 34.19 31.97 42.54 44.13 32.27 40.54
-F1 24.56 32.60 48.68 11.64 28.35 27.01 26.83 18.82 43.10 43.30 32.89 42.73
-HCI-V ACC 70.33 69.27 67.64 66.94 57.06 67.69 57.36 54.53 63.19 65.89 65.16 71.06
-F1 66.36 65.51 53.70 62.22 53.83 62.67 54.76 52.58 58.75 44.33 63.32 61.83
-HCI-A ACC 54.41 57.50 57.23 49.48 54.70 55.93 52.30 51.23 59.42 57.26 52.21 60.85
-F1 52.91 56.24 55.68 47.43 54.02 55.15 50.25 49.45 57.02 56.89 51.84 56.26
-HCI-VA ACC 31.56 28.30 44.92 35.07 34.84 30.09 26.99 27.21 41.54 37.43 43.02 36.92
-F1 26.29 27.58 21.65 27.19 27.96 24.71 21.95 20.92 38.08 19.60 36.07 32.99
-DEAP-V ACC 49.58 55.02 54.82 56.56 52.36 57.78 54.44 51.94 49.91 52.15 53.58 52.34
-F1 48.49 53.14 50.60 56.56 49.74 57.72 48.94 47.37 47.09 44.88 50.68 49.03
-DEAP-A ACC 51.48 50.99 26.48 41.08 48.94 49.73 45.90 44.09 49.91 43.86 50.05 50.59
-F1 50.95 50.09 25.39 41.05 48.94 49.17 45.56 41.51 47.09 40.46 47.79 46.38
-DEAP-VA ACC 24.58 27.84 16.06 17.31 25.41 23.90 24.64 20.89 25.66 19.09 30.92 20.49
-F1 24.70 24.36 15.19 16.87 24.44 22.58 23.24 15.16 24.95 13.02 30.98 18.41
+Method SVM CNN RNN GNN DNN Trans.
+EEGNet CDCN TSception ACRNN BiDANN R2G-STNN DGCNN GCBNet GCBNet _BLS RGNN NSAL-DGAT DBN DAN DANN PR-PL HSLT
+SEED ACC 75.08 58.81 68.23 64.01 49.71 78.37 78.88 82.55 80.56 76.64 76.55 78.62 71.88 60.35 62.00 76.54 64.83
+(19.73) (16.22) (20.35) (16.44) (13.15) (20.36) (17.34) (15.61) (16.98) (17.44) (16.92) (17.48) (19.02) (17.03) (21.95) (18.80) (20.47)
+F1 70.82 54.41 63.76 60.53 45.78 75.91 73.72 79.89 77.29 72.52 72.52 74.77 67.39 53.12 53.58 72.19 58.82
+(23.51) (17.59) (24.49) (18.51) (14.18) (24.60) (21.20) (18.93) (20.92) (21.20) (20.08) (20.95) (22.81) (19.70) (26.90) (22.85) (23.36)
+SEED-IV ACC 47.80 29.89 52.26 36.06 29.01 51.82 53.04 52.39 53.28 53.51 45.40 53.48 45.56 44.78 46.49 54.59 40.28
+(23.03) (13.53) (21.97) (15.12) (7.10) (23.71) (23.62) (24.32) (21.05) (22.45) (22.90) (21.21) (21.19) (23.84) (24.67) (22.13) (23.80)
+F1 40.17 26.59 45.26 32.77 19.80 44.65 46.48 45.94 46.26 46.91 38.24 47.09 37.61 37.18 37.76 48.24 30.92
+(21.68) (13.58) (23.00) (15.08) (5.42) (23.72) (24.76) (24.17) (22.27) (22.46) (23.09) (20.92) (20.68) (22.60) (23.89) (22.32) (24.47)
+HCI-V ACC 64.83 61.15 60.48 61.12 60.51 68.50 63.10 67.83 66.84 70.60 64.86 65.17 62.03 63.55 58.94 61.18 64.00
+(25.95) (16.76) (21.90) (15.52) (16.89) (24.18) (23.41) (22.40) (21.42) (21.05) (17.36) (20.52) (24.90) (22.19) (19.16) (12.19) (11.40)
+F1 55.99 50.35 51.73 50.51 49.39 58.74 53.30 54.78 54.61 58.96 50.41 58.85 52.84 48.68 44.96 41.16 55.77
+(27.80) (17.28) (23.11) (16.69) (15.70) (27.88) (25.38) (26.69) (25.08) (27.45) (20.34) (23.09) (27.04) (23.25) (21.05) (11.22) (13.12)
+HCI-A ACC 63.61 67.42 71.82 68.26 66.26 67.62 65.72 67.29 64.89 69.60 70.96 65.25 68.51 61.67 64.54 64.57 67.74
+(23.44) (21.71) (21.72) (23.10) (22.69) (24.58) (25.05) (27.73) (27.12) (22.09) (19.79) (26.37) (21.72) (25.08) (19.95) (23.95) (17.22)
+F1 50.99 54.50 62.89 56.29 55.17 57.77 54.82 58.04 57.43 57.78 57.66 53.40 57.18 48.25 50.16 50.50 58.42
+(24.89) (20.05) (24.99) (23.79) (23.20) (28.86) (24.52) (31.41) (29.50) (27.18) (25.39) (27.67) (26.63) (27.22) (23.81) (26.39) (19.50)
+HCI-VA ACC 46.29 38.32 52.00 40.00 41.00 51.31 48.41 53.06 49.15 49.24 49.46 53.35 44.38 41.88 40.68 44.27 46.99
+(28.42) (19.51) (26.05) (20.60) (21.58) (31.24) (25.36) (24.44) (26.16) (27.97) (23.20) (27.07) (26.78) (23.17) (22.11) (23.08) (20.76)
+F1 33.91 24.56 38.04 27.19 27.10 37.62 36.10 39.75 36.49 36.68 35.97 38.42 31.96 26.62 24.72 29.90 34.76
+(27.65) (13.71) (26.88) (13.65) (15.13) (27.81) (22.71) (26.01) (27.58) (27.25) (23.84) (27.48) (27.87) (17.51) (16.79) (20.68) (19.69)
+DEAP-V ACC 54.17 51.50 57.71 51.52 53.52 55.21 55.96 56.07 56.49 57.02 55.90 52.21 56.08 57.89 57.70 59.23 56.20
+(18.67) (11.57) (14.72) (9.54) (9.29) (19.70) (18.30) (17.15) (18.17) (15.07) (16.24) (16.19) (17.38) (19.18) (19.93) (14.69) (18.14)
+F1 49.73 47.85 53.41 47.33 48.31 47.23 49.84 49.08 50.36 51.40 47.25 48.87 48.61 50.97 50.23 47.42 48.21
+(18.92) (11.70) (15.50) (9.58) (7.77) (18.05) (17.77) (17.50) (19.57) (17.25) (17.55) (18.34) (19.33) (18.88) (20.97) (17.67) (18.73)
+DEAP-A ACC 63.49 61.30 63.37 57.49 61.83 63.57 63.78 62.68 65.95 61.07 66.09 60.31 64.60 60.51 62.38 62.15 59.74
+(16.72) (15.88) (14.18) (11.86) (14.32) (20.33) (18.92) (19.66) (17.61) (16.56) (13.91) (17.29) (19.42) (17.32) (16.78) (15.90) (18.82)
+F1 53.31 53.26 53.94 50.75 49.68 52.93 48.68 53.94 55.34 50.43 49.27 50.21 52.61 50.96 48.57 53.22 50.10
+(14.39) (13.05) (13.76) (11.30) (9.20) (20.26) (17.91) (20.10) (17.86) (16.36) (12.86) (18.91) (19.85) (16.53) (16.96) (17.99) (18.08)
+DEAP-VA ACC 37.32 39.41 38.08 35.68 38.20 39.30 40.75 41.86 38.80 37.51 44.53 38.78 39.50 38.07 41.71 42.58 43.34
+(17.24) (11.53) (15.52) (12.08) (11.34) (15.66) (15.79) (11.57) (16.23) (15.29) (14.35) (15.64) (13.99) (14.12) (15.51) (13.01) (14.49)
+F1 25.55 29.19 28.90 26.75 21.05 27.32 28.63 29.12 27.89 27.00 25.88 25.75 24.88 25.91 28.88 27.63 23.47
+(14.23) (10.21) (13.15) (8.30) (6.13) (13.23) (13.01) (10.92) (15.94) (13.52) (12.08) (11.92) (10.79) (10.98) (15.14) (11.43) (14.21)
+SEED-V ACC 34.88 36.02 65.48 31.45 34.10 49.17 47.21 57.59 60.68 63.71 55.31 52.37 28.64 22.37 27.77 49.75 47.17
+(18.42) (10.15) (11.49) (5.12) (7.21) (17.86) (19.22) (20.36) (18.42) (17.03) (12.88) (19.83) (15.62) (4.31) (17.64) (16.01) (16.82)
+F1 29.83 26.77 49.51 20.05 26.29 36.82 38.02 48.59 50.13 51.31 46.55 46.25 22.42 7.54 20.02 39.25 31.61
+(16.75) (8.11) (12.34) (2.59) (4.43) (14.18) (16.74) (20.88) (19.81) (18.91) (22.17) (19.34) (15.51) (2.16) (17.58) (15.51) (12.85)
+MPED ACC 21.49 18.43 20.89 19.01 18.88 25.75 24.43 22.93 22.83 22.65 20.72 22.19 22.22 19.47 21.24 20.40 18.47
+(7.33) (8.99) (5.48) (6.77) (9.97) (7.97) (9.28) (6.98) (8.30) (8.06) (5.34) (8.23) (5.17) (7.38) (7.52) (7.48) (4.28)
+F1 17.97 17.80 17.00 12.63 13.27 21.26 21.03 19.10 18.88 19.00 19.06 18.67 16.76 14.83 17.50 17.86 16.54
+(5.97) (6.58) (4.74) (6.80) (6.04) (6.81) (7.64) (6.28) (8.00) (6.70) (4.88) (7.45) (4.42) (7.39) (9.52) (6.61) (5.96)
+TABLE VI: The mean accuracies and F1 scores (and standard deviations) using the proposed benchmark for cross-subject EER experiment.
+The top two methods in each scenario are highlighted using bold and underlined formatting.
+Method SVM CNN RNN GNN DNN Trans.
+EEGNet CDCN TSception ACRNN BiDANN R2G-STNN DGCNN GCBNet GCBNet _BLS RGNN NSAL-DGAT DBN DAN DANN MS-MDA PR-PL HSLT
+SEED ACC 37.07 38.19 57.72 45.60 45.39 59.49 56.84 60.87 56.32 56.32 59.45 64.40 36.16 44.49 63.89 64.00 56.33 56.00
+(15.54) (17.25) (16.50) (8.71) (10.09) (15.92) (19.20) (15.62) (14.73) (9.91) (15.28) (19.00) (6.32) (7.68) (1.37) (18.75) (4.69) (14.76)
+F1 33.45 31.83 58.66 43.54 42.37 48.18 51.61 57.22 55.12 51.43 56.31 58.47 22.67 33.79 53.37 57.35 48.02 55.75
+(10.98) (12.33) (13.47) (9.72) (9.74) (18.23) (13.40) (16.41) (12.88) (6.54) (12.56) (19.50) (2.57) (5.60) (2.98) (18.67) (2.36) (14.73)
+SEED-IV ACC 28.98 28.19 31.03 34.19 31.97 42.28 42.39 42.54 32.27 40.54 48.23 52.29 36.82 38.24 25.55 56.07 43.42 30.33
+(4.09) (3.82) (9.45) (6.99) (7.02) (15.67) (16.02) (22.09) (11.43) (7.33) (16.14) (16.89) (12.81) (11.86) (2.54) (15.45) (5.48) (5.47)
+F1 24.56 28.35 27.01 26.83 18.82 37.31 42.76 43.10 32.89 42.73 45.43 45.39 32.60 26.04 11.99 48.68 42.37 11.64
+(5.31) (5.44) (6.63) (6.84) (6.80) (14.14) (17.54) (23.71) (12.53) (5.89) (14.89) (13.29) (12.81) (11.74) (1.96) (13.78) (2.39) (2.34)
+HCI-V ACC 70.33 57.06 67.69 57.36 54.53 73.92 64.76 63.19 65.16 71.06 64.94 64.04 69.27 61.50 63.61 67.64 68.79 66.94
+(8.74) (8.14) (8.15) (10.99) (9.87) (10.56) (13.28) (8.13) (6.76) (13.33) (6.72) (6.67) (8.09) (11.49) (6.57) (7.88) (10.43) (5.52)
+F1 66.36 53.83 62.67 54.76 52.58 66.02 59.54 58.75 63.32 61.83 52.50 38.96 65.51 50.97 38.80 53.70 54.90 62.22
+(5.89) (5.59) (10.57) (9.28) (4.22) (14.04) (13.41) (10.52) (2.33) (9.43) (12.87) (2.26) (12.54) (9.86) (2.22) (3.41) (14.21) (4.53)
+HCI-A ACC 54.41 54.70 55.93 52.30 51.23 49.98 56.21 59.42 52.21 60.85 57.36 51.00 57.50 56.42 49.06 57.23 55.48 49.48
+(14.26) (14.46) (19.96) (16.16) (16.66) (22.61) (13.33) (19.93) (10.83) (16.25) (22.72) (23.13) (9.23) (18.05) (23.43) (21.98) (16.73) (18.17)
+F1 52.91 54.02 55.15 50.25 49.45 48.53 55.24 57.02 51.84 56.26 56.90 32.41 56.24 45.35 34.68 55.68 54.45 47.43
+(11.20) (10.21) (11.60) (11.56) (13.70) (12.59) (10.27) (11.56) (12.32) (12.93) (11.78) (10.08) (5.51) (12.41) (10.36) (9.65) (9.92) (13.90)
+HCI-VA ACC 31.56 34.84 30.09 26.99 27.21 36.24 38.25 41.54 43.02 36.92 42.63 38.66 28.30 40.35 28.00 44.92 44.24 35.07
+(9.12) (9.32) (9.81) (8.81) (12.14) (19.56) (11.10) (9.82) (12.60) (18.54) (17.44) (22.77) (19.84) (22.83) (18.26) (20.67) (17.24) (13.21)
+F1 26.29 27.96 24.71 21.95 20.92 24.42 28.48 38.08 36.07 32.99 25.28 22.13 27.58 20.71 13.12 21.65 22.93 27.19
+(3.57) (3.12) (2.47) (6.24) (4.61) (5.91) (5.52) (2.43) (7.58) (12.41) (6.22) (8.92) (8.97) (7.13) (6.28) (8.94) (2.29) (5.64)
+DEAP-V ACC 49.58 52.36 57.78 54.44 51.94 53.49 49.97 49.91 53.58 52.34 52.65 52.22 55.02 51.88 44.32 54.82 51.77 56.56
+(6.77) (6.81) (5.00) (6.33) (7.44) (7.07) (7.19) (7.48) (8.72) (4.51) (8.09) (11.83) (7.19) (7.90) (5.94) (10.42) (10.60) (5.79)
+F1 48.49 49.74 57.72 48.94 47.37 45.96 47.75 47.09 50.68 49.03 50.22 45.10 53.14 44.61 30.61 50.60 40.21 56.56
+(6.34) (6.32) (6.00) (7.27) (6.66) (9.86) (7.75) (7.31) (11.20) (5.28) (6.23) (10.30) (7.16) (9.50) (2.58) (9.65) (13.48) (2.21)
+DEAP-A ACC 51.48 48.94 49.73 45.90 44.09 50.07 49.13 49.91 50.05 50.59 46.85 50.94 50.99 36.86 49.32 26.48 45.01 41.08
+(7.76) (7.72) (9.38) (8.16) (16.08) (9.01) (7.49) (9.27) (10.60) (19.22) (16.25) (9.60) (13.86) (10.64) (16.56) (13.21) (13.05) (14.32)
+F1 50.95 48.94 49.17 45.56 41.51 46.87 47.18 47.09 47.79 46.38 43.90 50.75 50.09 32.68 33.26 25.39 37.26 41.05
+(6.88) (6.75) (9.73) (7.29) (7.69) (8.20) (5.83) (6.12) (6.37) (11.85) (10.58) (8.46) (12.74) (9.55) (6.78) (10.91) (10.96) (13.94)
+DEAP-VA ACC 24.58 25.41 23.90 24.64 20.89 29.30 28.02 25.66 30.92 20.49 22.40 26.44 27.84 23.87 26.44 16.06 22.85 17.31
+(8.90) (8.86) (9.11) (8.69) (14.62) (8.34) (9.12) (8.22) (12.73) (11.13) (12.36) (14.74) (13.66) (6.08) (14.74) (14.23) (8.14) (14.77)
+F1 24.70 24.44 22.58 23.24 15.16 23.36 23.14 24.95 30.98 18.41 16.85 9.97 24.36 22.33 9.97 15.19 15.49 16.87
+(6.98) (6.49) (6.88) (7.46) (7.56) (4.24) (8.70) (5.32) (5.27) (10.33) (6.13) (4.58) (5.53) (5.47) (4.58) (5.02) (6.62) (5.17)
+SEED-V ACC 28.03 32.27 33.84 32.63 27.03 36.58 37.23 36.25 30.84 35.68 36.45 41.56 32.18 20.82 23.99 44.33 33.77 30.59
+(10.91) (3.57) (9.87) (4.90) (12.76) (12.89) (12.47) (2.61) (6.51) (13.22) (5.65) (13.19) (6.56) (5.45) (3.98) (12.09) (2.24) (12.13)
+F1 19.54 22.22 27.45 25.33 18.27 32.72 31.48 29.77 27.70 30.02 31.93 35.33 25.29 14.97 17.96 38.28 29.64 26.54
+(11.93) (3.42) (11.00) (5.61) (11.65) (12.71) (13.62) (2.25) (3.64) (16.43) (10.42) (12.84) (7.71) (4.54) (4.08) (13.48) (3.00) (5.87)
+MPED ACC 19.56 20.33 17.83 17.65 17.53 25.75 22.29 19.64 18.82 19.01 18.85 19.07 17.83 19.10 17.77 20.67 18.76 17.94
+(3.29) (5.44) (1.09) (1.14) (2.41) (7.97) (5.40) (5.46) (2.69) (1.63) (3.35) (0.57) (1.09) (1.89) (1.44) (2.15) (0.39) (4.21)
+F1 16.36 17.27 15.31 15.65 15.28 21.26 18.19 16.25 15.42 15.42 15.22 15.55 15.31 17.17 15.54 16.79 16.10 15.27
+(3.75) (3.60) (1.19) (1.70) (4.11) (6.81) (5.44) (5.56) (2.89) (2.07) (2.37) (1.48) (1.19) (1.89) (1.50) (3.51) (0.60) (3.92)
+(a) subject-dependent
+(b) cross-subject
+Figure 5: The model performance scoring on subject-dependent and cross-subject tasks based on ACC and F1.
 V-A Model Reproduction
-To ensure the reliability of our benchmark, we conducted a meticulous replication of the representative EER methods using the original configurations and compared them with the results reported in the original paper.
-Table IV presents both our reimplementation outcomes and those reported in the original study with the gap.
-Additionally, we detail the specific original experimental setups, clearly illustrating the differences in experimental details among various models.
-All reproduced models have been integrated into our LibEER algorithm library, allowing for direct usage.
+To ensure the reliability of our benchmark, we conducted a systematic reproduction of the representative EER methods by strictly following the experimental settings reported in their original papers, including model configurations and dataset usage.
+Table IV presents both our reproduction results and the original reported results, along with the performance gap.
+Considering that SEED and DEAP are the two most widely used public datasets, and all included methods have been evaluated on at least one of them, we reproduced the results specifically on SEED and DEAP.
+If the gap was within 10%, we regarded the reproduced implementation as sufficiently reliable for subsequent unified benchmark comparisons.
+In addition, Table IV provides the specific original experimental setups used in the reproduction experiments, clarifying the differences in settings across different models compared to our unified benchmark.
+All verified models have been integrated into our LibEER algorithm library for direct use.
 From the results in the Table IV and the replication process, we can draw the following conclusions:
-(1) The reproduced model performances are generally slightly lower than the results reported in the original paper, averaging 2.95% lower, indicating a satisfactory overall replication quality.
-Notably, the reproduction results for RGNN, DBN, R2G-STNN, GCBNet_BLS, and ACRNN showed more significant declines compared to the original paper.
-Conversely, some scenarios in MS-MDA, as well as CDCN and HSLT, exhibited improvements in replication results, while the performances of other models were comparable to those in the original study.
+(1) The reproduced model performances are generally slightly lower than the results reported in the original paper, averaging 3.85% lower, with a maximum difference of less than 10%, indicating a satisfactory overall replication quality.
+The reproduction results for ACRNN, R2G-STNN, GCBNet_BLS, RGNN, DBN, and DAN showed more significant declines compared to the original papers ($>$5%).
+The reproduction results for BiDANN, DGCNN, GCBNet, NSAL-DGAT, and PR-PL showed smaller deviations ($<$5%).
+Conversely, some scenarios in CDCN, TSception, MS-MDA, and HSLT exhibited improvements in the reproduced results.
 The primary reason for these discrepancies is the lack of detailed descriptions regarding model specifications and experimental setups in some papers.
@@
 Consequently, even when attempts are made to replicate the results as closely as possible to the descriptions in the original work, it remains challenging to achieve identical outcomes.
-(2) There are significant differences in the experimental setups among various methods.
-The paper employs different datasets for each model, utilizes various data preprocessing techniques, executes distinct experimental tasks, and applies different data partitioning and performance evaluation methods.
-The substantial variability in experimental configurations prevents fair comparisons among the models.
-This underscores the necessity of establishing a benchmark under a unified and rational experimental framework.
+(2) There are substantial inconsistencies in the experimental setups reported in the original papers, including differences in datasets, preprocessing, tasks, and evaluation protocols.
+Notably, all original settings do not adopt a three-way split (train/validation/test), which increases the risk of overfitting and inflating performance due to the lack of validation-based model selection.
+This substantial variability in experimental configurations prevents fair comparisons among the models.
+Therefore, we established a unified benchmark under a consistent and rational experimental framework, as described in Section III, where all methods are evaluated using standardized preprocessing, dataset usage, a three-way split, and consistent evaluation metrics.
 V-B Model Comparison
 As shown in Table IV, the implementation details of different methods vary significantly, making it difficult to fairly compare their performance and draw reliable conclusions for the advancement of the EER field.
-To this end, we conduct a fair comparison of ten selected methods based on the unified and reasonable benchmark proposed in Section III.
-The results for the two experimental tasks are reported in Tables V and VI, respectively.
+To this end, we conduct a fair comparison of seventeen selected methods based on the unified and reasonable benchmark proposed in Section III.
+The results for the two experimental tasks are reported in Tables V and VI, which clearly present results that differ significantly from those in Table IV.
 We select SVM [36], one of the most commonly used machine learning algorithms, as the baseline model.
@@
 From these results, we make the following key observations.
-Observation 1. The graph-structured information between EEG signal channels is crucial for EER tasks. Several GNN-based models achieved superior performance.
-To provide a clearer illustration of the performance of each method under the benchmark, we ranked each method’s performance in each scenario.
-The highest-ranking method among the $n$ methods receives $n$ points, while the lowest-ranking method receives 1 point.
-The total score for each method across all scenarios is then accumulated, and the scoring results for both experimental tasks are presented in Fig.
+Observation 1.
+Modeling the spatiotemporal information of EEG signals is crucial for EER tasks.
+Several GNN-based and RNN-based models achieved superior performance.
+In addition, integrating transfer learning can effectively alleviate inter-subject variability in EEG data, demonstrating considerable potential.
+Figure 6: Radar chart of model performance across different datasets on subject-dependent and cross-subject tasks.
+To provide a clearer illustration of the performance of each method under the benchmark, we designed a scoring scheme that simultaneously considers accuracy and F1-score across all datasets.
+Specifically, for ten datasets (SEED, SEED-IV, HCI-V, HCI-A, HCI-VA, DEAP-V, DEAP-A, DEAP-VA, SEED-V, MPED), all $n$ methods are evaluated based on both accuracy and F1 score, yielding a total of 20 rankings (10 datasets × 2 metrics).
+In each ranking, the top-performing method receives $n$ points, while the lowest receives 1 point.
+The final score of each method is computed as the sum of its points across all 20 rankings, as shown in Fig.
 5 (a) and (b).
 From Fig.
-7 (a), it can be observed that in the subject-dependent task, DGCNN achieved the best performance with a total score of 140.
-Following closely are GCBNet and its variant GCBNet_BLS, both scoring 132 points.
-CDCN and RGNN ranked fourth and fifth with scores of 126 and 105, respectively.
-In contrast, EEGNet, Tsception, and ACRNN performed relatively poorly.
+5 (a), it can be observed that in the subject-dependent task, DGCNN achieved the best performance with a total score of 275.
+Following closely are GCBNet and its variant GCBNet_BLS, scoring 259 and 255 points, respectively.
+CDCN and BiDANN ranked fourth and fifth with scores of 235 and 234, respectively.
+In contrast, EEGNet, Tsception, ACRNN, DAN, and DANN performed relatively poorly.
 In the cross-subject task (see Fig.
 5 (b)), the overall performance ranking of each model is similar to that in the subject-dependent scenario.
-Notably, the DBN model, which has a relatively simple architecture, rises to third place in the cross-subject scenario.
-In contrast, the MS-MDA model, specifically designed for cross-subject scenarios, shows significant variability across different datasets, resulting in a moderate ranking in the overall score.
+Several GNN-based methods still achieved relatively superior performance.
+BiDANN and R2G-STNN, two models based on RNN and transfer learning that already performed well in subject-dependent settings, achieved even better performance in the cross-subject scenario, ranking second and third, respectively.
+EEGNet and DBN obtained better results on this task compared to their performance in subject-dependent settings.
+The MS-MDA model, specifically designed for cross-subject scenarios, resulted in a moderate ranking in the overall score.
 In terms of model types, the four GNN-based methods demonstrated superior performance on both tasks, providing strong evidence that the graph-structured information between EEG signal channels is crucial for EER tasks.
 In contrast, the CNN-based methods exhibited varying levels of performance on both tasks.
-The DNN and Transformer-based methods showed average results on the subject-dependent task, while the DNN-based methods performed better on the cross-subject task.
-The RNN-based methods performed relatively poorly.
-However, for the Transformer and RNN-based methods, as only one model was selected from each category, the results may contain some occasionality.
-Based on the current results, we recommend that researchers in the EER field focus more on using GNN-based models when designing their models and emphasize the channel graph structure information inherent in EEG signals.
-(a) subject-dependent
-(b) cross-subject
-Figure 7: The Model Performance Scoring on subject-dependent and cross-subject tasks.
-(a) subject-dependent
-(b) cross-subject
-Figure 8: Radar chart of model performance across different datasets on subject-dependent and cross-subject tasks.
+The DNN and Transformer-based methods showed relatively weaker overall performance, though some individual models demonstrated promising results on specific tasks.
+However, for the Transformer-based methods, as only one model was selected, the results may contain some occasionality.
+RNN-based models exhibited notable performance divergence, with both BiDANN and R2G-STNN ranking within the top five and even top three across both tasks, while ACRNN consistently ranked last.
+This may be attributed to the incorporation of transfer learning in BiDANN and R2G-STNN, which effectively models inter-sample variability.
+However, the two DNN-based transfer learning models, DAN and DANN, performed poorly, indicating that beyond employing transfer learning, it is equally important to capture the intrinsic spatiotemporal characteristics of EEG signals through architectures such as RNNs and GNNs.
+Based on the current results, we recommend that researchers in the EER field prioritize models capable of capturing the spatiotemporal dynamics of EEG signals, such as those based on RNN or GNN, and consider incorporating transfer learning techniques to enhance model generalizability across subjects.
+(a) Subject-dependent task on the SEED-IV dataset
+(b) Cross-subject task on the SEED-IV dataset
+(c) Subject-dependent task on the HCI-VA dataset
+(d) Cross-subject task on the HCI-VA dataset
+Figure 7: Experimental results on sample length for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
+(a) Subject-dependent task on the SEED-IV dataset
+(b) Cross-subject task on the SEED-IV dataset
+(c) Subject-dependent task on the HCI-VA dataset
+(d) Cross-subject task on the HCI-VA dataset
+Figure 8: Experimental results on feature type for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
 Observation 2.
@@
 Consequently, further in-depth investigation is necessary to adequately confront the challenges posed by the substantial differences in EEG data among various subjects.
-Observation 3: The scarcity of EEG data limits the effective representation learning capabilities of deep learning methods, leading to some deep learning approaches performing even worse than traditional machine learning methods, such as SVM.
-As demonstrated in Tables V and VI, several carefully designed deep learning algorithms did not outperform the baseline SVM, as reported in their original studies.
-A significant contributing factor to this discrepancy is the scarcity of EEG data, characterized by both limited quantity and scale of datasets.
-Deep networks typically necessitate substantial data volumes to fully leverage their potential for representation learning.
-Additionally, many studies lack validation sets in their experimental designs, which can easily result in overfitting and inflated performance metrics in contexts with limited data.
-In contrast, our rigorous experimental setup has highlighted these shortcomings.
-Therefore, a prudent strategy would be to utilize various data augmentation techniques [44, 43, 42] to expand the dataset and to implement strategies aimed at reducing model overfitting.
-Observation 4.
-The data collection methods for different EEG datasets vary, resulting in differences in data quality and task difficulty.
-Several methods performed better on the SEED dataset, which has a more reasonable data collection approach.
-To visually illustrate the performance differences of various models across different datasets, we created radar charts for both experimental tasks, as shown in Fig.
-8.
-From the results presented in Fig.
-8, we summarize that models performed better on the three-class SEED dataset compared to the binary classification tasks in DEAP and HCI.
-We believe this is due to the higher label quality in SEED, where discrete emotion labels are easier for self-assessment, while continuous VAD labels tend to be more subjective.
-Additionally, SEED provides expectations for the type of emotions elicited by stimuli, whereas DEAP and HCI do not.
-However, as the number of labels increases, all models demonstrate similar performance on the four-class tasks of SEED-IV, DEAP-VA, and HCI-VA, with relatively poorer performance observed in DEAP-VA.
-(a) Subject-dependent task on the SEED-IV dataset
-(b) Cross-subject task on the SEED-IV dataset
-(c) Subject-dependent task on the HCI-VA dataset
-(d) Cross-subject task on the HCI-VA dataset
-Figure 9: Ablation study results on sample length for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
-(a) Subject-dependent task on the SEED-IV dataset
-(b) Cross-subject task on the SEED-IV dataset
-(c) Subject-dependent task on the HCI-VA dataset
-(d) Cross-subject task on the HCI-VA dataset
-Figure 10: Ablation study results on feature type for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
-TABLE VII: Ablation study results on training methods for subject-dependent task on the SEED-IV and HCI-VA datasets.
+TABLE VII: Experimental results on training methods for subject-dependent task on the SEED-IV and HCI-VA datasets.
 Method SEED-IV HCI-VA
@@
 (19.55) (19.65) (19.00) (23.12) (20.77) (17.40)
-TABLE VIII: Ablation study results on training methods for cross-subject task on the SEED-IV and HCI-VA datasets.
+TABLE VIII: Experimental results on training methods for cross-subject task on the SEED-IV and HCI-VA datasets.
 Method SEED-IV HCI-VA
@@
 F1 26.63 25.18 23.42 27.70 25.07 18.76
+Observation 3: The scarcity of EEG data limits the effective representation learning capabilities of deep learning methods, leading to some deep learning approaches performing even worse than traditional machine learning methods, such as SVM.
+As demonstrated in Tables V and VI, several carefully designed deep learning algorithms did not outperform the baseline SVM, as reported in their original studies.
+A significant contributing factor to this discrepancy is the scarcity of EEG data, characterized by both limited quantity and scale of datasets.
+Deep networks typically necessitate substantial data volumes to fully leverage their potential for representation learning.
+Additionally, many studies lack validation sets in their experimental designs, which can easily result in overfitting and inflated performance metrics in contexts with limited data.
+In contrast, our rigorous experimental setup has highlighted these shortcomings.
+Therefore, a prudent strategy would be to utilize various data augmentation techniques [51, 50, 49] to expand the dataset and to implement strategies aimed at reducing model overfitting.
+Observation 4.
+The data collection methods for different EEG datasets vary, resulting in differences in data quality and task difficulty.
+Several methods performed better on the SEED series datasets, which feature a more reasonable data collection approach.
+To visually illustrate the performance differences of various models across different datasets, we created radar charts for both experimental tasks with ACC and F1 score, as shown in Fig.
+6.
+From the results, we summarize that models performed better on the three-class SEED dataset compared to the binary classification tasks in DEAP and HCI.
+We believe this is due to the higher label quality in SEED, where discrete emotion labels are easier for self-assessment, while continuous Valence-Arousal-Dominance labels tend to be more subjective.
+Additionally, SEED provides expectations for the type of emotions elicited by stimuli, whereas DEAP and HCI do not.
+As the number of labels increases, various models achieve relatively similar average accuracy on SEED-IV, HCI-VA, and SEED-V, with only DEAP-VA showing notably lower performance.
+However, the performance differences among models are more pronounced on SEED-IV and SEED-V compared to the other two datasets.
+Moreover, due to the more balanced label distribution in the SEED series datasets, their F1 scores are significantly higher than those of the other datasets.
+The MPED dataset, which contains the largest number of emotion categories, poses significant challenges to most models, indicating that fine-grained recognition of human emotions in the EER field still requires further investigation.
+(a) Subject-dependent task on the SEED-IV dataset
+(b) Cross-subject task on the SEED-IV dataset
+(c) Subject-dependent task on the HCI-VA dataset
+(d) Cross-subject task on the HCI-VA dataset
+Figure 9: Experimental results on random seed for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
 V-C Discussion of Experiment Settings
-In this section, we select multiple representative methods from each kind of EER method and conduct ablation studies on three crucial factors, i.e., sample length, feature type, and evaluation methods.
+In this section, we select multiple representative methods from each kind of EER method and conduct comparative experiments on four crucial factors, i.e., sample length, feature type, evaluation methods and random seeds.
 To be specific, SEED-IV and MAHNOB-HCI with four classification labels are taken as the dataset and SVM, DBN, HSLT, CDCN, ACRNN, and DGCNN will be selected as the representative methods.
@@
 To explore the effects and impact of different sample lengths, we examined lengths of 1s, 3s, 5s, 7s, and 9s, with the results presented in Fig.
-9.
+7.
 As illustrated in Fig.
-9, in the subject-dependent scenario, the performance of most methods tends to decline as sample length increases.
+7, in the subject-dependent scenario, the performance of most methods tends to decline as sample length increases.
 We attribute this trend to the sparsity of data in the subject-dependent context; shorter sample lengths yield a greater total number of samples, which enhances the learning capacity of deep models.
@@
 V-C2 Data preprocessing Feature
-In the ablation study on feature types, we utilized both DE and PSD features, along with their respective variants processed by Locality Differential Sparse (LDS), as inputs.
+In the comparative experiments on feature types, we utilized both DE and PSD features, along with their respective variants processed by Locality Differential Sparse (LDS), as inputs.
 The results are presented in Fig.
-10.
+8.
 As illustrated in Fig.
-10, in the subject-dependent scenario, the DE and PSD features demonstrate comparable performance, while both feature types processed by LDS show significant improvements.
+8, in the subject-dependent scenario, the DE and PSD features demonstrate comparable performance, while both feature types processed by LDS show significant improvements.
 In the cross-subject scenario, the DE feature exhibited the lowest performance, whereas the other three features displayed similar efficacy.
-This finding underscores the utility of LDS as a valuable feature processing technique, suggesting that the selection of specific features should be adapted flexibly according to the experimental context and the unique characteristics of the model.
+This finding highlights LDS as a valuable feature processing technique, suggesting that the selection of specific features should be adapted flexibly according to the experimental context and the unique characteristics of the model.
 V-C3 Evaluation Methods
-In the ablation study on model training methods, we aimed to explore how different training approaches impact model performance.
-We compared three training methods: (1) the method used in our benchmark, which selects the epoch with the best performance on the validation set (as “Ours” in Tables VII and VIII); (2) selecting the model’s result from the last epoch (as “R1” in Tables VII and VIII), where, for a fair comparison, the model’s training set includes the validation set; and (3) selecting the epoch when the loss value on the validation set decreases by less than 1% within the first 10 epochs (as “R2” in Tables VII and VIII).
+In the comparative experiments on model training methods, we explored how different training approaches impact model performance.
+We compared three training methods: (1) the method used in our benchmark, which selects the epoch with the best performance on the validation set (as “Ours” in Tables VII and VIII); (2) selecting the model’s result from the last epoch (as “R1” in Tables VII and VIII), where, for a fair comparison, the model’s training set includes the validation set; and (3) selecting the epoch when the loss on the validation set decreases by less than 1% within the first 10 epochs (as “R2” in Tables VII and VIII).
 Tables VII and VIII show that the training method used in our benchmark achieved the best performance in the majority of cases and obtained the highest average performance.
@@
 These findings highlight the importance of using a reasonable and appropriate training method for the effective training of EER models.
+V-C4 Random Seeds
+The random seed determines the reproducibility of experimental results, yet it is often unspecified in existing studies.
+To investigate the impact of random seed settings on model performance, we conducted comparisons using seeds set to 2023, 2024, 2025, 2026, and 2027.
+The results are shown in Figure 9.
+Two key conclusions can be drawn from the results.
+First, the random seed has a significant impact on experimental outcomes, with different models exhibiting considerable performance variation under different seeds.
+This is because changing the random seed affects both the random data splitting process and the stochastic behaviors during model training, such as parameter initialization, mini-batch ordering, and dropout, all of which can lead to variations in the final results.
+Among all the methods, CDCN is most affected by the random seed, with performance differences reaching approximately 20%.
+In contrast, HSLT and ACRNN are relatively less affected, though their variations still reach around 10%.
+Second, in the majority of cases, these models exhibit similar variation patterns when the random seed changes.
+This suggests that using a unified random seed across all models can help ensure the fairness of comparisons to some extent.
+This finding underscores the importance of adopting the same random seed for all methods in the EER field and explicitly reporting the seed value in publications to ensure fairness and reproducibility of results.
 VI Future Direction
@@
 Experimental Settings.
-LibEER includes several key experimental setups to facilitate research, including four widely recognized public datasets, common data preprocessing techniques, two main tasks, common data partitioning methods, and reasonable performance evaluation methods.
+LibEER includes several key experimental setups to facilitate research, including six widely recognized public datasets, common data preprocessing techniques, two main tasks, common data partitioning methods, and reasonable performance evaluation methods.
 However, LibEER does not cover all experimental configurations used in the EER field.
-Some studies use other datasets like DREAMER[47], Amigos[48], and MPED[49].
+Some studies use other datasets like DREAMER[54] and Amigos[55].
 Others develop unique preprocessing methods, including customized filtering, artifact removal[11, 58, 59], denoising[56, 57], normalization techniques[57, 11], and extracting other types of features [15, 52, 18].
@@
 By standardizing the experimental settings, including dataset usage, data preprocessing methods, experimental task selection, data splitting strategies, and evaluation methods, LibEER addresses the prevalent issues of inconsistency and reproducibility in the EER field.
-Through its unified PyTorch codebase and the comprehensive implementation of over ten representative models, our benchmark ensures fair comparisons and accurate evaluations of models in both subject-dependent and cross-subject scenarios.
-Additionally, through fair comparison experiments and ablation studies, we present findings and discussions that aim to inspire researchers in the field.
+Through its unified PyTorch codebase and the comprehensive implementation of seventeen representative models, our benchmark ensures fair comparisons and accurate evaluations of models in both subject-dependent and cross-subject scenarios.
+Additionally, through fair comparison experiments and investigation of key experimental settings, we present findings and discussions that aim to inspire researchers in the field.
 We believe that LibEER will not only facilitate the development and benchmarking of new EER models but also promote standardization and transparency, thereby advancing research and applications in this domain.